This page: http://monitor.mikes.fi/ionclock/

Notes - Comb - Ion - Cavity - DAQ

Archived notes: index_2019 index_2020
Machines and IPs:
Strontium/linux .45.99 (strontium.ad.vtt.fi)
win .45.23 (ESPTE1M012-ioni.ad.vtt.fi)
kavi .45.190 (kavi.ad.vtt.fi)

News and notable results

  • 2021-04-30: trapping again after IR-camera BBR-characterization. New machined alu C-beam mount
  • 2021-01-25: first ions trapped after x-mas break. Intensifier+PCO needed for beam-alignment
  • 2021-01-22: new mu-shields delivered. inner and mid-shields installed and fit ok
  • 2020-12-03: Initial measurement of repumper-shift: -2600 (C1) to -2900 Hz (C4)
  • 2020-11-18: New Ion page for clock-run data
  • 2020-11-04: clock run with new HelRes (now 14.424 MHz) and new B-beam direction, but only 800us probe
  • 2020-09-24: Move 2019 results to index_2019
  • 2020-09-24: Secular frequencies with tickler + photon-correlation
  • 2020-09-23: new 14.404 MHz Magic-RF Helical-Resonator
  • 2020-07-02: Fourier limited 4 Hz linewidth with 180ms probe-pulse. Night, no metro.
  • 2020-06-11: Second clock run, OD2 with 4ms probe, only C1, C2 peaks.
  • 2020-05-17: D5/2 lifetime, random-telegraph method with 405nm and fit
  • 2020-04-24: First clock run! 2.5ms probe, stability 2e-13/sqrt(tau_s)
  • 2020-04-01: trapping single ion again, after 405 laser broke
  • 2020-01-31: Move comb-plots to comb
  • 2020-01-15: For temperatures, see DAQ
  • 2019-12-19: Micromotion map, 3D pictures
  • 2019-11-29: Cavity mode number, Comb mode number. Cavity absolute frequency.
  • 2019-11-08: -C2 Zeeman component linewidth estimate of 800 Hz.
  • 2019-11-05: Zeeman peaks with B-shield (1.6 uT estimate).
  • 2019-11-04: Now with first B-shield. Laminated plywood + metglas.
  • 2019-07-12: Zeeman peaks in ambient (66uT estimate) B-field. First clock-spectrum!
  • 2019-07-10: Initial clock-transition search...
  • 2019-07-10: Page created.
  • 2019-06-18: ion photo and secular frequencies

  • 2021-05-28 Friday

    11:50Z Switch lasers off.
    11:00Z Switch lasers on. No ion in the trap?

    2021-05-26 Wednesday

    14:28Z turn off lasers. leave ion trapped

    Detuning-scans at V12cm=0V:


    NOTE 0.5*axial plotted, to get all datapoints close to eachother

    Summary of results:
    [-5.8 x 26.6721 x] V -> ( 958830, 1003244, 1975363) Hz
    [-5.0 x 27.7201 x] V -> ( 982384, 1025684, 1950869) Hz
    [-4.0 x 29.0301 x] V -> (1010912, 1053126, 1919859) Hz
    [-3.0 x 30.3401 x] V -> (1038747, 1079756, 1888188) Hz
    [-2.0 x 31.6501 x] V -> (1065803, 1105771, 1855910) Hz
    [-1.0 x 32.9601 x] V -> (1092152, 1131241, 1823320) Hz
    [ 0.0 x 34.2701 x] V -> (1117918, 1156046, 1789792) Hz
    V12diff = 0.2451 V constant
    V34diff = 50.0363 V constant
    

    shift per volt = (+25800, +24800, -33500)
    axial scan 1788-1792 kHz
    rhi scan 1154-1158 kHz
    rlo scan 1116-1120 kHz

    uMotion A 2.6%, B 5.5%, C 7.2%
    Set Vbias = [0 x 34.2701 x]

    shift per volt = (+25500, -32600)
    axial scan 1821-1825 kHz
    rhi scan 1130-1134 kHz

    uMotion A 3.4%, B 4.1%, C 7.1%
    Set Vbias = [-1 x 32.9601 x]

    shift per volt = (+2700, +26000, -32300)
    axial scan 1854-1858 kHz
    rhi scan 1104-1108 kHz
    rlo scan 1064-1068 kHz

    uMotion A 3.9%, B 2.4%, C 6.9%
    Set Vbias = [-2 x 31.6591 x] V

    shift per volt = (+27800, +26600, -31700)
    axial scan 1887-1891 kHz
    rhi scan 1078-1082 kHz
    rlo scan 1037-1041 kHz

    uMotion A 3.4%, B 3.2%, C 7.3%
    Set Vbias = [-3.0 x 30.3401 x]

    Observed shift per volt = (+28500, +27400, -31000) Hz
    axial scan 1918-1922 kHz
    rhi scan 1051-1055 kHz
    rlo scan 1009-1013 kHz

    uMotion A 4.4%, B 4.6%, C 7.6%
    set Vbias = [-4 x 29.0301 x] (x not changed)

    Observed shift per volt V12cm (x, y, z) = (+29200, +27700, -31300) Hz
    axial scan 1947-1954 kHz
    rhi scan 1023-1027 kHz - could use even more tickler than 7dB att
    rlo scan 980-984 kHz


    uMotion A 4.4%, B 5.1%, C 7%
    rlo ca rhi ca +2.7 kHz/0.1V, expect +22 kHz shift
    axial ca -3 kHz/0.1V, so expect -24 kHz shift
    Set new Vbias = [-5.0 0.2451, 27.7201, 50.0363]

    (x, y, z) = (958830, 1003244, 1975363 Hz ) Yesterday: (959049, 1003521, 1975940)

    fit fails - too wide scan?

    repeat rhi

    Vbias = [-5.8 0.2451 26.6721 50.0363]

    Detuning scans.
    NOTE: figure P422 and P1092 not measured, just guesses. Program doesn't include window-transmission difference for A/B and C beams


    uMotion 100kphot: A 4.7% in 13.9s, B 3.7% in 18.9s, C 7.5% in 13.8s

    Fit uses -1 MHz correction, so 2*208 - 436 - 1 = -21 MHz (this agrees with the scanned points in the figure)
    Is 208 MHz a good set-point for cooling?

    ion appears maybe 20s after completed detuning-scan! (only points -23 MHz from 208MHz were scanned, no points in blue direction)
    try detuning-scan to get cold ion?
    no ion during first 60s with only A-beam, or during about 60s with ABC beams
    detuning is 208 MHz
    11:06Z turn on lasers.

    2021-05-25 Tuesday

    synced with gitlab, new secular_prediction.py program

    14:25Z lasers off. leave ion trapped

    New version 'v2' of detuning-scan program with approximate lineshape fit.
    TODO: automatically measure background-counts before or after the scan (or take this from blue-detuned data points?). Add parameters for P422_A,B,C and P1092


    uMotion A 6.1% in 10s, B 3.3% in 15.8s, C 4.3% in 10.5s

    lasers on, ion still trapped

    2021-05-24 Monday

    14:31Z turn off lasers. leave ion trapped.

    Test how repeatable axial secular results are. RID14875 - RID14884 (10 repeats). each measurement ca 3 minutes.
    (measurements via remote desktop, no activity in the lab)
    f0 = [1975659, 1975684, 1975645, 1975599, 1975633, 1975598, 1975663, 1975607, 1975562, 1975502]
    FWHM = [671, 559, 580, 610, 695, 628, 508, 630, 795, 657]
    

    Standard deviation for resonance frequency ca 50Hz, for FWHM ca 75 Hz.
    Systematic effects due to lab temperature change or helres-drift, cooling-laser intensity/polarization are probably (much) larger?



    Trap parameter results:
    Vbias -> [omega_secular x, y, z] (uMotion A, B, C)
    [-5.8 0.2451 26.6721 50.0363] -> [ 959049 1003521 1975940] (7.7%, 1.2%, 2.8%)
    [-5.9 0.2451 26.5411 50.0363] -> [ 956205 1000753 1978978] (7.1%, 1.6%, 2.5%)
    [-6.0 0.2451 26.4101 50.0363] -> [ 953220  997869 1982072] (not measured)
    (lunch)
    [-6.0 0.2451 26.4101 50.0363] -> [ 952917  997615 1981450] (7.9%, 2.1%, 3.5%)
    
    seculars at V12cm = -5.8 V

    get uMotion A 7.7%, B 1.2%, C 2.8% Go to Vbias = [-5.8 0.2451 26.6721 50.0363]

    Seculars at V12cm = -5.9 V

    get uMotion A 7.1%, B 1.6%, C 2.5%
    Go to Vbias = [-5.9 0.2451 26.5411 50.0363] (change is +0.1 * dVbias)

    Trap parameter analysis suggests voltage-vector that doesn't change field at ion:
    dVbias = [1, 0, 1.31, -0.03] (V34diff change is small - ignore?)

    Repeat secular measurement:
    953.2 / 997.9 / 1982.1 kHz @ Vbias = [-6 0.2451 26.4101 50.0363], uMotion
    (these are +315 PPM / +301 PPM / +339 PPM compared to values measured before lunch )
    Ion went dark during axial scan!? (too much tickler with 20+10 dB setting!) Fluorescence recovered with detuning-scan


    Lunch

    RID14835 C-beam, axial secular mode with C-beam, 20+20 dB
    RID14836 C-beam, rhi 20+10 dB scan 995-1000 kHz (peak height 15%, could use less attenuation?)
    RID14838 C-beam rlo 20+10 dB scan 950-955 kHz, 50 points (side-to-side blur at resonance in camera image, with good imagination..)
    952.9 / 997.6 / 1981.4 kHz


    PMT counts with 1000s gate, time-series ca 60s
    RID14825 B-beam
    RID14827 A-beam
    RID14829 C-beam

    feels like there is quite a lot of variation in P422
    P1092 8.90 mW
    422 beam power (signal-background): A 0.55-0.08 uW, B 0.90-0.15 uW, C 0.79-0.07 uW


    RID14821 C-beam scan, max 21.7 kcounts, C fluorescence 8859 (1384) / 167 (16)
    RID14818 B-beam detuning scan. max 12.7 kcounts, B fluorescence 5531 (228) / 64 (8) bright/dark
    RID14815 A-beam detuning scan. max counts ca 21k/1000ms. A fluorescence 9342 (195) / 65 (9) bright/dark.

    ID14808 detuning-scan, A-beam, 1000ms gate, -23...+10 MHz from 208 MHz set-point in 60 steps
    A-fluorescence in 1000ms is 6854 (240) / 57.3 (8), repeat 5938 (281) / 67 (6)

    uMotion A 7.9% in 15.9s, B 2.1% in 17.5s, C 3.5% in 14.5s

    This ion loaded Monday 2021-05-17, and spent several nights + the weekend un-cooled.
    07:57 ion appeared! after re-locking to Rb-line!? (ionization lasers were on also)
    no ion in the trap? (tried 405nm and 461nm lasers also) tried 422 detuning-scan.
    07:41Z turn on lasers

    Notes on geometry and voltages

    Current voltages are from uMotion optimization run 2021-05-18 ID14607 NPL search, [ 0.24512619 26.41013719 50.03631515 ], V12cm = -6 V

    2021-05-21 Friday

    13:38Z: Switch off lasers for the weekend. Leave ion in the trap.

    uMotions: A: 7%, B: 3%, C: 4%

    12:13Z: Switch lasers on. Ion still in the trap!

    2021-05-20 Thursday

    PMT off.
    15:42Z ion is still stored and cooled. Turn off lasers, leave RF on.


    P1092=11mW assumed. Note new C-beam focus seems OK, almost as good as A. Matlab-code fix for C-beam (IR window transmission)

    422-beam power before chamber: A 0.65-0.08 = 0.57uW, B 1.15-0.18 uW = 0.97 uW, C 0.9-0.15 = 0.75 uW

    RID14790 C-beam detuning scan
    RID14787 C-beam fluorescence 10743 / 220 bright/dark (note >2-fold more dark-counts! 422-scatters into PMT!?)

    RID14787 B-beam detuning scan, 1000ms gate
    RID14785 B-beam 1000ms fluorescence 7785 / 67 bright/dark

    RID14783 detuning-scan with A-beam, 1000ms gate
    RID14784 A-beam 1000ms fluorescence @ 208 MHz is 9742 / 71 bright/dark (compare to ca 90 counts for the most blue-detuned points in the scan

    axial around 1.982 MHz (same as before)


    uMotion A 7.5% in 13.8s, B 1.7% in 17.7s, C 4.5% in 13.5 s

    last experiments ca 17 local on Wednesday, so ion stored ca 23 hours without cooling
    ion appears on camera ca 6-7 seconds after opening the 422-shutter. Detuning 208 MHz as usual
    12:55Z turn on lasers
    Mid-shield was removed y-day, inner shield is installed

    2021-05-19 Wednesday

    Set RF back to 10 Vpp and switch off lasers

    RF 4 Vpp, Monitors: CH3: 0.720000 V, CH4 1.400000 V :

    Photon correlation tickler:
    Lower radial 20+12 dB:


    Upper radial 20+14 dB:

    Axial 20+29 dB:

    Coarse:
    Axial 0.7637 MHz


    Radials 0.34 MHz and ~0.448 MHz (??)


    RF 7 Vpp, Monitors: CH3: 1.200000 V, CH4 2.419999 V :
    uMotions: A: 10%, B: 2%, C: 5%

    Photon correlation tickler:
    Lower radial, 20+10 dB:

    Upper radial, 20+12 dB:

    Axial, 20+25 dB:


    Coarse tickler, 0+11 dB:

    C-beam:
    Radials, 0.6517 MHz, 0.7146 MHz:

    Axial: 1.363 MHz



    RF 10Vpp, Monitors: CH3: 1.719999 V, CH4 3.479998 V :

    Photon correlation ticker:
    Lower radial, 20+8 dB:

    Upper radial, 20+10 dB:

    Axial, 20+23 dB:

    Coarse tickler, 0+8 dB:

    A-beam:
    ID14673 Radials, 0.9529 MHz, 0.99753 MHz:


    ID14674 Axial 1.982 MHz


    Optimize beams

    uMotions: A: 4%, B: 4%, C: 3%

    Remove mid-shield and install 2nd RF-monitor probe

    10:18Z Switch lasers on. Ion still in the trap!

    2021-05-18 Tuesday

    14:36Z turn off lasers. leave ion trapped.

    Install fixed 20dB attenuator. Axial with 20+23 dB

    Try axial mode without fixed attenuator, 0+30 dB dds attenuation.
    Too much tickler amplitude.


    15:53+03 uMotion 100kphot: A 3.6% in 13.2s, B 2.7% in 20.9s, C 3.9% in 22.2s

    14640, 14641, 14642 improve program so beam-letter is included in figure? (include also photons/point?)
    Upper radial:

    Lower radial: (easier to excite, so peak 2-fold higher with same tickler-attenuation)
    height also inversely proportional to FWHM (cooling rate) ?


    Upper radial just below 998 kHz (harder to exite wrt. lower radial)
    A, 0+23 dB. FWHM is large, so mode aligned along beam??

    Lower radial around 952.8 kHz
    Correlation with 0+23 dB attenuation
    C, 0+23 dB. Peak height 0.6, too high?
    B, 0+23 dB.
    A, 0+23 dB.

    RID14625 photon-correlation tickler, 20+20 dB, A-beam. 20kphotons/pt. inner+mid shield installed. ASI camera
    A, 20+20 dB
    A, 20+23 dB
    B, 20+23 dB, FWHM is almost 2-fold larger than A
    C, 20+23 dB,

    Switch to ZWO ASI camera

    uMotions at new optimum: A: 6%, B: 1%, C: 2%
    ID14607: NPL search, [ 0.24512619 26.41013719 50.03631515]

    Intensifier powered from Siglent lab-supply; axial:


    Ion reappeared!

    Ion was lost?
    ID14591: NPL search for uMotion minimum. 422 lost lock, fail.

    uMotions: A: 8%, B: 20%, C: 22% at [-6, 0.239, 25.919, 51]

    06:00Z Ion still in the trap!

    2021-05-17 Monday

    Tried with another +12V source. Similar result. Investigate intensifier power box?

    After removing intensifier power supply:


    ID14563: Switch intensifier off:


    ID14562: After installing mid-shield, now 0+28 dB:

    Install mid-shield.

    ID14555: Intensifier on

    ID14552: After installing shield, intensifier still off.

    Install second half of inner shield.

    ID14551: Intensifier off, second peak diminishes slowly?



    Switching off intensifier, still two peaks visible? Relaxation time?
    ID14543: Removed foil, now two peaks visible for C beam axial:


    ID14510-13: Intensifier switched off, lower radial, photon timestamp correlation, 20+11 dB




    ID14508: Intensifier: lower radial, photon timestamp correlation, 20+11 dB

    ID14507: Intensifier: lower radial, 0+8 dB

    Install intensifier and set to max gain.

    ID14494: ZWO ASI: lower radial, 0+12 dB


    Results below were obtained with ZWO ASI camera

    ID14490: Lower radial, C beam, photon correlation, 20+11 dB

    ID14489: Upper radial, C beam, photon correlation, 20+11 dB

    ID14488: Axial, C beam, photon correlation, 20+25 dB:


    ID14487: Axial, C beam, attenuation 0+25 dB. 1.9815 MHz
    ID14485: Upper radial, C beam, attenuation 0+8 dB. 0.9972 MHz
    ID14484: Lower radial, C beam, attenuation 0+8 dB. 0.95228 MHz
    ID14482: Coarse tickler: Radials at approx. 0.95234 MHz and 0.9972 MHz

    uMotions at V34diff = 51 V: A: 4%, B: 24%, C: 24%

    Sweeping V34diff gives new guess for minimum for A: 51 V.

    uMotions: A: 60%, B: 67%, C: 70%

    11:20Z Try to load. I = 1.6 A. Ion! at 9min 30s

    Check 461 dispenser. Set I from -196.83 mA -197.11 mA

    10:49Z Try to load. I = 1.6 A. No ion at 10 min. Flush trap and continue. After 6 min two ions appear. Flush trap and continue. No ion for 10 mins.

    Check resonator Z-matching: Minimum 1.01 at 14.4236 MHz. 1.05 at 14.424 MHz. Adjust minimum to 14.424 MHz, now 1.01.

    10:29Z No ion in the trap.

    2021-05-12 Wednesday


    14:13Z turn off lasers

    uMotion A 45%, B 41%, C 46%

    axial with photon-correlation, B-beam, 20+25 dB attenuation, 10 kphotons/pt.


    Upper radial with correlation method. 20+11 dB attenuation, B-beam, 10kphotons/pt


    coarse tickler 0+3 dB. half of axial (?) and upper radial


    coarse tickler 0+14 dB, lower radial, B-beam

    photon-correlation method, 20+11 dB attenuation, B-beam, 10 kphotons/pt - looks ok!?


    lower radial with 0+14 dB, ca 950.7 kHz


    0+14 dB artiq DDS, scan around upper radial

    0+4 dB artq dds tickler shows both of these.
    Upper radial scan: peak below 0.99 looks like half the axial mode, around 0.996 is the upper radial

    Switch to ASI USB3 camera, to get more magnification

    siglent siggen 1.8 Vpp hiZ amplitude for upper radial, manual tickler
    upper radial

    0.996 and 0.95 MHz radial modes. Upper harder to excite.
    Axial mode 0+24/26 dB attenuation

    siglent siggen 280mVpp in hiZ mode - for manual tickler
    coarse tickler shows resonances at 1.979 and 1.980 MHz

    moving camera up/down and l/r increased PMT counts to >200/10ms (from 20/10ms)

    12:36Z ion! (tried adjusting 461nm laser-current during loading)
    12:14:15Z load with 1.60A
    1s camera exposure

    detuning to 185 MHz for loading
    throw away two ions
    align A-beam (not much improvement in fluorescence counts 20-40/10ms)

    At 10Vpp see axial mode around 1.980 1.982 MHz

    reducing trap drive to 4Vpp shows two ions
    detuning to 208 MHz
    11:39:15Z (ca 10mins)
    11:29Z load with 1.60A (on 6.5 loading took up to 12 mins!?)
    Bias voltages from previous settings -6 0.239 25.919 55 V
    Detuning 185 MHz for loading
    RF on, 10 Vpp drive, Rigol monitor voltage 3.46 Vpp
    old camera with intensifier

    11:27Z turn on lasers.

    Try to record images of secular modes, to see what is causing double-peaked secular resonances?

    2021-05-10 Monday

    NOTE: results below are due to filter series-resistance together with 10 MOhm input impedance of DMM!

    Sanity-check of bias voltages?
    V1 on bias-voltage source dispplay: -6.23901 V
    V1 from bias-voltage BNC-output (no cable connected): -6.2379x V (last digit x unstable, Keysight 34461A DMM)
    V1 from end of pink SMA-cable -6.2380x V (not connected to filter-board)
    V1 from green screw-connector (with blue wire to trap conencted): -6.2116x V (voltage is 0.0264 V more compared to unloaded bias source? has this been measured before?)

    V2 on bias-source display: -5.76101 V ( 0.5*(V2-V1)=0.239 V is the V12diff programmed - seems OK with V12cm=-6V programmed)
    V2 from bias-voltage BNC-output: -5.7601x V V2 from end of pink SMA-cable (not connected to filter-board): -5.7602x V
    V2 from green screw connector -5.7356x V (0.0245 V more than from unloaded bias-source)

    2021-05-06 Thursday

    14:40Z Turn off lasers, leave ion trapped

    back to -6V V12cm
    Test with V1 and V2 bias-box outputs disconnected and V1/V2 cables terminated with 50R. Still double-peak. Now 10Vpp drive
    1.783 and 1.7848 MHz

    two axial modes ca 2 kHz apart seen also with siglent generator connected to piece of wire, held in front of the IR-window (IR-side inner shield removed)

    Axial secular scanning with Siglent generator (no ext ref). No external attenuator. -12 dBm output (into 50 ohm). Weaker excitation ~1.559 MHz, stronger excitation ~1.5609 MHz.

    RID14286 Try correaltion method with more photons 40 kphot/pt. C-beam 20+20 dB att.


    8 Vpp C-beam axial, 20+9 dB att, 80 ms gate.
    8 Vpp C-beam axial, 20+9 dB att, 40 ms gate. longer gate reduces false peak? measurement is long enough to show 422 power variations
    8 Vpp C-beam axial, 20+9 dB att, 20 ms gate
    8 Vpp C-beam axial, 20+9 dB att, 10 ms gate
    ID14277 try C-beam correlation method, with a lot more attenuation 20+20 dB. 10 kphotons/pt
    8 Vpp C-beam axial, 20+20 dB att
    8 Vpp C-beam axial, 20+15 dB att

    8 Vpp C-beam axial, 20+9 dB att
    8 Vpp B-beam axial, 20+9 dB att (B should have more projection onto z?)
    8 Vpp A-beam axial, 20+9 dB att
    8 Vpp A-beam axial, 20+8 dB att
    8 Vpp A-beam axial, 20+7 dB att
    uMotion A 3.5%, B 5.7%, C 8.9%
    Set RF-drive 8Vpp

    9Vpp A-beam axial, 20+7 dB att. NOTE sudden shift down -2kHz(?) when working close to the trap
    9Vpp C-beam axial, 20+7 dB att
    9Vpp B-beam axial, 20+7 dB att
    9Vpp A-beam axial, 20+7 dB att
    9Vpp A-beam axial, 20+5 dB att

    uMotion A 4.3%, B 4%, C 8.5%
    Set RF-drive 9 Vpp

    shift of about -0.5 kHz during a few minutes?
    A-beam axial, 20+5 dB att
    C-beam axial, 20+5 dB att
    B-beam axial, 20+5 dB att

    Shape of Rb-lock signal changes a lot when moving close to the laser...
    Rb-lock unlocked...

    Axial ca 1.9745 MHz, 20+4 dB with all beams
    axial, 20+4 dB att
    uMotion A 6%, B 2.4%, C 6%

    Ion seems to be still trapped, fluorescence re-appears a few s after opening 422-shutter
    07:23Z turn on lasers.

    2021-05-05 Wednesday

    Turn off lasers, intensifier. Leave RF on and ion trapped

    Double-peak is some nonlinear effect due to too strong tickler??

    axial, 20+7 dB att
    axial: lower resonance seems weaker, upper is strong with 28 dB att.
    axial, 0+28 dB att
    axial, 0+25 dB att
    upper radial 0+8 dB att
    lower radial 0+12 dB att
    coarse tickler 0+8 dB shows radials ca 944 and 989 kHz. upper harder to excite.


    back to V34dif=55V: get uMotion A 2.6%, B 5.9%, C 7.5%
    uMotion A 68%, B 65%, C 60%
    All beams illuminate ion ok
    detuning to 208 MHz
    13:40:30 signal - why 12 min loading-time?
    13:28:30Z load with 1.60A
    Detuning to 185 MHz
    RF on/off to throw away the bright/dark pair

    one axial mode maybe 947 kHz
    sometimes shifted sideways position when scanning axial tickler
    Ion jumping side-to-side seems to indicate one bright ion and one dark ion?
    Axial mode ca 1.979 MHz. Coarse tickler with 0+28 dB seems too strong

    uMotion A 8%, B 68%, C 44%
    A-optimum at V34diff 50V
    Trap voltage down to 4Vpp - no clear sign of 2 ions.
    uMotion 100kphot: A 66%, B 77%, C 6% (but 11% 2nd harmonic)

    A-alignment OK, B some optimization needed, C completely dark - optimization needed
    set detuning 208 MHz
    12:59 signal! weak cloud (not cooled?) at first - then cold and small image (later this looks like it was one dark and one bright ion)
    12:53Z load with 1.60A.
    405nm power ca 3.5mW
    rf off/on to empty trap

    12:47 no signal... check Rb-lock. check 461-dispenser 196.90 mA (close to previous setting)
    12:34:30Z load with 1.60A, 1.58W initially. Detuning ca 186 MHz. Yesterday loading time was ca 4 min
    Back to voltages -6.0 0.239 25.919 55 V
    Install new filter-board - ion was lost during start of removal of old filter-board

    Set 0 0 33.719 55 V (uMotion A 23%, B 49%, C 34%
    11:03Z turn off monitor voltage logging with rigol

    At 0 0.239 33.719 55 V get uMotion: A 0.6%, B 4.8%, C 5.5% - seems to confirm 1:1.3 ratio for V12cm:V34cm
    Previously 1.3-fold change in V34cm gave micromotion-null. 25.919+1.3*6 = 33.719 V
    Down to V12cm=0V in 0.5V steps. This gives uMotion A 70.4%, B 64%, C 67%
    Voltages were -6 0.239 25.919 55 V
    Try to lower abs(V12cm) before changing filter-board

    Fluorescence appears again some 10s of seconds after re-locking to Rb-line.
    Rb-signal changes shape a lot, and is unstable. Now quite a lot of negative DC-offset (previously positive)
    10:51 cooling laser Rb-lock unlocked!?
    (after lunch)

    Something in RF-drive/helres not very stable, now axial suddenly higher again, and monitor voltage also higher?
    briefly seem to see 1.87 MHz or higher axial, then goes lower to below 1.86 in a few minutes!?


    Axial mode with correlation method
    B-beam 0+30 dB, 20 kphotons/pt
    B-beam 0+30 dB, 20 kphotons/pt, wider range 1858-1863 kHz. cooling doesn't work for upper resonance!
    B-beam 20+13 dB, 20 kphotons/pt, (added physical attenuator)

    Check spectrum of tickler-DDS with Siglent SA. Looks OK, no visible sidebands or spurs.
    With 18 dB DDS attenuation ca -8 dBm peak with spur-free SNR of ca 80 dB

    correlation method for upper radial
    B-beam 0+20 dB
    B-beam 0+18 dB, 935-945 kHz

    correlation method for lower radial. no physical attenuator. C-beam, 10kphotons/point

    20kphotons/pt:
    C-beam 20kphotons/pt, 0+16dB attenuation
    A-beam 20kphotons/pt, 0+16dB attenuation
    B-beam 20kphotons/pt, 0+16dB attenuation. ion not cooled at lower resonance
    B-beam 20kphotons/pt, 0+20dB attenuation. ion image/cooling ok through resonance


    axial secular now around 1.86 MHz (down from 1.886 MHz at end of Tuesday)
    Radials ca 892 khz and 940 kHz
    Upper radial is harder to excite.
    C-beam 0+27dB, 10ms gate
    C-beam 0+8dB, 10ms gate
    C-beam 0+10dB, 10ms gate


    Detuning is 208 MHz
    Fluorescence counts in 10ms: A 158/4, B 111/4, C 129/6
    uMotion 100kphotons: A 5% in 9.3s, B 3% in 12.3s, C 4.5% in 10.5s

    07:15Z lasers on, ion seems to be still stored. camera image maybe a bit defocused?
    This is maybe the 2nd (?) time ion has been stored without cooling over night?
    sometime during the night, monitor-voltage dropped from 3.36Vpp to 3.2Vpp (rigol scope)

    2021-05-04 Tuesday

    14:30Z turn off lasers. leave rf on and ion trapped

    B-beam counts 66/0.22 in 10ms
    B-beam power ca 0.6 uW
    C-beam 0+28dB, 100ms gate
    C-beam 0+30dB, 100ms gate
    B-beam 0+28dB, 100ms gate
    C-beam power ca 0.6 uW
    C-beam ca 70/1.5 in 10ms
    misalign 422 fiber-coupling (laser breadboard)

    C-beam power ca 1.3 uW (too much?). Fluorescence counts in 10ms is 137/7
    uMotion 100kphot: A 1.9%, B 4.6%, C 6.2%
    Back to 208 MHz
    0+26dB, 100ms gate
    0+26dB, 100ms gate

    Try detuning 185 MHz
    0+26dB, 50ms gate
    0+26dB, 100ms gate (note different rage)

    Try changing V12cm +0.2V to -6.0, and V34cm +1.3*0.2V to 25.919V - should change axial/radial ratio?
    0+8
    0+10 dB
    axial 0+24 dB
    axial 0+27 dB
    axial 0+26 dB
    axial 0+26 dB (high to low)
    axial 0+26 dB (lo to high)

    B-beam, sweep over both radial modes
    0+10 dB
    0+8 dB now also upper radial?

    Try to measure axial mode around 1.7-1.9 MHz? 0+22dB attenuation

    (sweep now from high to low)
    (sweep now from high to low)
    (low to high)
    ca 1893 and 1895 kHz.
    radials were 898 and 899 kHz

    A-beam, 0+10dB attenuation 0.9-1 MHz, 2000 pts, 10ms gate (down from 100ms previously). check how repeatable results are?
    ID13969, ID13970



    Zoom in on 0.88-0.92 MHz and repeat: now 0+8 dB attenuation


    0+10 dB
    0+12 dB
    0+12 dB

    peaks seem to move around a lot?? changing Rf-amplitude?
    further coarse-tickler scanning 0+8 dB shows peaks: ca 862 kHz, 867 kHz, and 944 kHz !?

    ID13938: Maybe two peaks? Too much power?
    Attempt to measure with photon correlation.

    ID13933: Two-peaked lower radial now with 13 dB instead of 3 dB!
    Disconnected new tickler coupling (to lower electrode). Attached old coupling capacitor to upper electrode terminal.

    ID13912 photon-correlation tickler 933-935 kHz, 0+29dB att, 100pts, B-beam 10 kphot/point

    ID19311: Radial (lower) sweep (coarse): 0.93236 and 0.93372 MHz
    ID13903: Axial secular sweep (coarse): 1.9632 and 1.96496 MHz
    ID13902: Radial (upper) secular sweep (coarse): 0.978 and 0.980 MHz

    setting V34diff to 55V gives A 2%, B 5%, C 2.5%
    lowering V34diff to 56V gives A 18%, B 22%, C 21%

    uMotion 100kphot A 39.9%, B 42%, C 42% (1st harmonic looks ok, with only little 2nd harmonic)
    set detuning to 208 MHz - image now much brighter
    10:32:10 signal!
    10:28Z load with 1.60A. 1.56W initially, ca 1.60W after 1 min, 1.607W at 2 min,
    set detuning 185 MHz

    no cooling during lunch -> ions seem to have disappeared?
    when lowering trap-drive from 10Vpp down the glitch is between 3.6->3.5 Vpp
    two ions visible when lowering trap-drive to 4 Vpp. blink-test with 405nm confirms this
    uMotion A 15%, B 3%, C 3% (but 11% 2nd harmonic?)
    set detuning 208 MHz
    08:50 ion signal!
    08:43 load again with 1.60A. Now with better 461nm current value
    08:41 signal - blinking? looks like many ions, so throw away
    08:27 load with 1.60 A dispenser, 461 current driver at too high set-point. Set I=-197.0mA at 10 min.
    08:15Z turn on lasers

    2021-05-03 Monday

    no time for trapping today, turn off lasers
    no sign of ion...
    08:15Z turn on lasers, check if ion is still stored?

    2021-04-30


    Turn off lasers. leave RF on and dark ion trapped (until Monday?)

    422 power check after detuning scans (compare to before values below): A 1.0 uW, B 1.6 uW, C 1.3 uW
    Detuning-scans with improved beam alignment now looks better:
    ID13862, A-beam, 1s gate. Fluorescence at 208MHz in 1s gate: 13462 / 64. scan is -23 MHz to +12 MHz from nominal detuning in 60 steps
    ID13865 B-beam 1s gate. Fluorescence(208MHz, 1s): 12106 / 55
    ID13868 C-beam (new!) 1s gate. Fluorescence(208MHz, 1s): 10208 / 386 (compare the dark backround count to blue-detuned values in the scan?)

    uMotion now A 9%, B 3.6%, C 3.9%
    aligning ASE-beam improves counts from 50/10ms up to >200counts/10ms - a 4-fold improvement.
    422 beam power check: A 0.7 uW, B 1.3 uW, C 1.3 uW - similar as before.

    uMotion check: A 4.9%, B 3.2%, C 3.2% - these look OK, but why are detuning-curves so bad??
    Typical detuning curve:


    wide scan with 0+6 dB:

    With 0+6 dB attenuation also a double-peak around 933 kHz!??

    Stronger tickler possibly reveals lower radial mode around 933 kHz
    radial sweep, 0+17 dB attenuation. 978.3 kHz peak, 980 kHz peak (half of axial?)

    tickler-spectrum with 0+28 dB attenuation:


    ID13810 detuning-scan C-beam, 100ms gate. Fluorescence at 208 MHz in 100ms is 326.53/14.98 (note almost 2-fold more dark-counts with C-beam, 422-beam scatters more into camera/PMT?)
    ID13807 detuning-scan B-beam, 100ms gate. Fluorescence at 208 MHz in 100ms is 285.1/6.46
    ID13803 detuning-scan A-beam, 100ms gate, -23 and +12 MHz from 208 MHz. Fluorescence at 208MHz in 100ms is 355.6/7.9

    Fluorescence levels (10ms) are A 35.6/0.86, B 27.8/0.72, C 35/1.24
    Algorithm suggests 0.239 25.659 57.290 where we get: (100kphot) A 6.4% in 31s, B 0.6% in 36s, C 1.7% in 32s
    ID13788 uMotion minimization 2x2x2 grid.
    (Note NPL minimization program previously had wrong V12cm=-5V setting, now set to -6.2V as used while manually scanning)

    (Note V12cm = -6.2 V is held constant) Based on this 0.202 25.5 57.5 should be A+B optimal? Get A 11%, B 2%, C 8% ( by luck? we found also a C-optimal point!?)
    V34cm 26V: A-opt (26, 57), B-opt (26, 58)
    V34cm 27V: A-opt (27,56) and B-opt (27, 59)
    Try same scans at V34cm=28V: B-opt (28, 60), A-opt (28, 54)
    further scanning shows B-optimal point ca (V34cm, V34diff) = (29.301, 61) V . should now change something so A-optimum at (29.301, 52.5) coincides?
    now A 1.7%, B 82%, C 76% at voltages 0.202 29.301 52.5
    scanning V34 diff shows A-minimum at 52.5V

    at this point uMotion A 48%, B 74%, C 68%
    optimum found 0.202 29.301 55.057
    ID13755 uMotion scan 4x4x4 grid, 50 kphot/point, estimate duration 50min (lunch)

    uMotion A 8%, B 70%, C 72%
    scanning V34diff shows A-min at 61V
    uMotion A 64%, B 81%, C 77% (2nd harmonic content lower)
    08:37:20 ion signal!
    08:28 try to load with 1.60A dispenser current. Rb-lock of cooling laser seems a bit jumpy, very sensitive to air-currents above laser (shield this with alu-plate or plexi?)
    trap RF OFF and the ON, to empty trap
    Ions seem to be lost
    08:10Z turn on lasers.
    two ions were left in the trap yesterday, lasers turned off.

    2021-04-29

    Blink test doesn't look great. two or more ions


    Suggested optimum 0.199 23.38 65.08 - here get A 23% B 22% C 27%
    ID13727 try uMotion scan 3x3x3 grid with 10kphot/point
    Try detuning 208 MHz as usual?

    suggested optimum 0.253 26.89 59.51. At this point A 33%, B 23%, C 21% (all with a lot of 2nd harmonic) - make sure we have only one ion!?
    ID13719 try uMotion minimization with 2x2x2 grid, 50 kphot, ranges [0.2, 0.25], [26, 27], [58, 59]

    Now uMotion 100kphot A 5.5%, B 36%, C 29%
    Try scanning V34diff with A-beam: A-contrast minimum at ca 58.4 V
    based on bias-voltage program we are now at -6.2 0.239 26.3 62.4 V
    uMotion A 55% B 19% C 21% - but large 2nd-harmonic correlation, so bias voltages require adjustment
    Note detuning now ca 190 MHz (because cooling didn't work previously with 208 MHz)
    Install IR-side inner shield. C-beam passes through the shield hole OKish
    C-beam gives ca 50 counts/10ms (maybe focus is a bit off)
    Install new C-beam. Requires 0.5" extension lens-tube, so that fiber-connector doesn't interfere with Polaris M4 mounting screws

    NOTE: 1.958/2 = 0.979 - so one peak in the radial spectrum might be half the axial mode?
    With 0+26 dB attenuation axial mode around 1.958 MHz

    Tickler with 0+13 dB attenuation shows two peaks at 0.978 and 0.979 MHz (don't know if this is the higher or lower radial mode)


    uMotion A 17%, B 15%
    mostly side-to-side adjustment of B-beam gives 70 counts/10ms for B also.
    ca 70 counts/10ms from A-beam, very little from B-beam
    detuning to 208 MHz, ionization lasers off.
    CMI FLIR IR-camera monitors dispenser, and window/chamber black tape reference areas
    11:00Z: Try to load, 1.7 A, Ion! after 7 min.

    10:48Z: Try to load, 1.6 A. No ion after 10 min.

    7:38 UTC: start working in the lab.

    2021-04-27

    Check 461 dispenser: -196.95 mA, 4.462 kOhm.+

    2021-04-08

    turn off lasers - not sure why loading didn't work?? :(
    12:31Z no signal... dispenser off
    no signal in 8-9 min, try lowering RF-drive to 4.0 Vpp (same as earlier today when loading)
    12:07Z load with 1.60A dispenser (1.54W initially... rising to 1.604W)
    empty trap by turning RF off, and then back on
    set detuning to 185 MHz

    seems ion was lost. not sure why - higher RF-voltage might give unstable trapping parameters?

    ion went dark!?
    Upper radial see two peaks with 0+15 dB: 1.45822 and 1.458975 MHz

    Upper radial seems easier to excite?
    Matlab secular predicition: 1.436, 1.469 3.06 MHz
    Monitors now 5.2 and 2.16 Vpp
    Set Keysight to 4.8 Vpp - 20% higher than before

    Monitors 4.4 Vpp and 1.76-1.8 Vpp
    lower radial 1.179495 MHz with 0+2dB
    lower radial ca 1.18 MHz. Found with 0+2dB attenuation.
    radial 0+10dB, 1.218889 MHz
    axial again 2.490206 MHz, 20+9 dB attenuation

    axial at 2.4894 MHz, 0+24 dB attenuation

    with 0+6dB get asymmetric peak (too much heating)
    Coarse tickler 0+10dB att finds one peak at 1.2184 MHz

    Lunch-break...

    Counts at 208 MHz are 43.8/1.3 in 10ms
    Also try this with the lights off in the prep-room: (3-4/10ms lower background in 10ms, so 30-40 lower in 100ms?)

    Repeated detuning scan with A-beam


    optimizing PMT/pinhole-position gives slightly better counts 50/10ms
    at -6.2 0.239 26.3 62.4 detuning-scan with A-beam shows a better slope over the detuning-range
    Counts still quite low 40/5 in 10ms
    (detuning scan with 100ms gate)


    V34d 62.4V slightly better giving A 1.7%
    at -6.2 0.239 26.3 63V get uMotion A 11% and B 31%
    Further scanning V34diff shows A-minimum around V43dif=63V
    NOTE: IR-side inner shield not present. Imaging this side with the FLIR camera

    A-beam PMT counts 42/6 in 10ms (lights on in the prep-room give higher dark-counts)

    B-beam detuning scan.


    A-beam detuning scan. For secular correlation we want a large slope at 208 MHz detuning - not so great here..
    Scan range (from 208MHz) is -23 MHz to +12 MHz


    leave at V34diff=45V. uMotion A 50%, B 49%
    Try V34diff 44..48V - no clear minimum. Try V34diff 48..52V - optimum might be at a lower voltage
    Try sweeping V34diff 52...56 V and search for B-minimum? no clear minimum
    uMotion 100kphot: A 83% in 28s, B 82% in 36s
    PMT counts B-beam 34.5/1.5 in 10ms (similar to A, so maybe this is due to micromotion)
    PMT counts with A-beam are 39.9/1.8 in 10ms - seems a little low (micromotion, or alignment?)
    Set detuning to 208 MHz - giving much brighter image

    Both A and B beams give fluorescence - so their alignment is ok
    dispenser off, photoionization off.
    07:50:40 signal!
    07:45 try to load with 1.60A (1.54W initially)

    intensifier + PCO camera (200 ms exposure)in use. Objective attached and camera focus to "7" on the micrometer-translator. Field of view should be large, almost see top/bottom inner electrodes in image
    Latest bias voltages in ARTIQ are -6.2, 0.239, 26.3, 54 V (these probably in use now). Actual voltages on device screen -6.439 -5.961 -27.7 +80.3 V
    RF on at 4.0 Vpp, giving monitors 4.4 Vpp and 1.8 Vpp with Rigol scope
    Cooling laser on. Rb-signal seems to wander a lot? Optimize AOM-angle for 185 MHz detuning

    2021-04-07

    Keysight siggen glitches: 4.0->4.1 Vpp when moving up, 3.6-3.5 Vpp when moving down

    2021-04-01 Thursday

    Setting RF to max (and testing) at 11:03 UTC

    Setting RF drive to nominal 2.9 Vpp at 10:36 UTC with amplifier
    Setting RF drive to nominal 10 Vpp at 10:30 UTC without amplifier

    Adjusted impedance matching of resonator. Now optimal SWR = 1.01 at 14.424 MHz

    Setting RF drive to nominal 10 Vpp at 10:14 UTC without amplifier

    Setting RF drive to nominal 2.9 Vpp at 09:26 UTC with amplifier

    Using trap drive of 1.5 Vpp:

    ID13539: axial: B-beam: 28.5 dB, peaks at 0.8874 and 0.89087


    ID13553: radial: B-beam: 0+2.5 dB, peaks at 0.39152 and 0.388136


    ID13522: radial: B-beam: 0+1 dB, peaks at 0.4431156 and 0.445226


    ID13516: radial: B-beam: 0+15 dB, peak at 0.48653



    Using trap drive of 2.5 Vpp:
    ID13506: axial: B-beam: 0+28.5 dB, peaks at 1.4898 and 1.4922


    ID13494: radial: B-beam: 0+13 dB, peaks at 0.764045 and 0.766067


    ID13490: radial: B-beam, 0+2.5 dB, peaks at 0.744638 and 0.745797


    ID13479: radial: B-beam, 0+3 dB, peaks at: 0.7044925 and 0.70652.


    Three peaks ??? :



    2021-03-31 Wednesday

    Leave cooling laser and ASEs on for the night. Switch off intensifier and PMT.

    Switched back to 208 MHz detuning

    Bias voltages: [-6.2, 0.239, 26.3, 54]

    ID13382: B-beam detuning scan


    Secular: correlation method with 20 dB attenuator:
    ID13446: Axial 20+17 dB attenuation: split into 2 peaks approx. 2 kHz apart!


    Seculars: old method: no additional attenuation:
    ID13454: axial: B-beam, 0+27 dB, with 185 MHz detuning.


    ID13445: axial: B-beam, 0+28 dB, peak at 1.737692


    ID13423: lower radial, B-beam 0+3 dB att. peaks at 0.826577 and 0.8282684


    ID13360: upper radial, B-beam, 0+13dB att. 0.8785 MHz

    ID13380: Radial scan, peaks at 0.877826 and 0.880029


    V34diff: 54 V, V34common: 26.3: A: 2%, B: 2%

    55.6 V V34diff: A: 70%, B: 1%

    ID13332: Trying lower radial with B :

    Set voltages to: -6.2, 0.239, 28.0873, 52

    49 V V34diff: A: 54%, B: 81%
    Search for rough umotion minimum: 52V V34diff: A: 8%, B: 67%

    Install inner shield ASE side (both parts of inner B shield now installed).

    Ion from Monday still in the trap!

    RF drive is now 2.9 Vpp, with RF amp and 35 dB input attenuation.
    Switch to Rigol scope for RF monitoring.

    2021-03-29 Monday

    ID13299: Trying lower radial with B :

    ID13298: Trying axial with B, two peaks? :

    A 63%, B 3% at V34diff = 69 V

    A 4%, B 62% at V34diff = 65 V

    A 32%, B 42 % (no C-beam installed)
    maybe we have too much micromotion?
    axial secular seems to drift higher as measurement proceeds?
    12:58Z ID13266 correlation-measurement of axial, 1.724 - 1.73 MHz, 20 + 17 dB attenuation

    maybe around 1.727 MHz - not so great shaped resonance..
    coarse tickler 0+20 dB att axial 1.8-2.0 MHz -> somewhere in 1.72-1.74 MHz range
    upper radial ca 885.4 kHz, 0+10 dB tickler att.
    coarse tickler shows lower secular 833.4 kHz.

    A-beam gives ca 100counts/10ms
    blink-test with 405nm (low power) seems to show only one ion.
    disconnect 422nm fiber to use it for aligning the PMT-splitter. Fiber-end at PMT-position, watch for centered 422-spot with the IR-side camera
    Currently no PMT-splitter installed...
    coarse tickler-sweep 800 kHz - 1 MHz seems to show two resonances - OK
    found A-beam alignment also (at first only B-beam gave fluorescence)
    12:03 signal! from B-beam? dispenser off
    11:58:30 load with Dispenser 1.7A
    RF signal-gen 2.9Vpp, monitors 3.12 Vpp and 1.33 Vpp

    A-beam power:
    461nm 132uW
    422nm A 0.7 uW, B 1.0 uW
    405nm 3.5 mW
    repumper 11.4 mW
    clearout 62 uW

    Check 461 dispenser: -197.13 mA, R 4.462

    align intensifier + pco camera to inner electrodes
    connect 2nd monitor to scope. monitors 4.11 and 1.68 Vpp
    remove IR-side inner shield
    remove outer shield, remove mid-shield
    08:13Z lights on in the lab, remove finnfoam-boards
    (note summer time now, UTC+3)

    2021-03-26 Friday

    13:36Z dispenser off - could not get any cloud or signal visible??
    12:59Z dispenser 1.66A
    lab dark, intensifier to 5V, ca 400-800 ms exposure on PCO camera
    422nm detuning 185 MHz
    RF generator 2.8 Vpp (+attenuation + RF PA), monitor 3.1Vpp and 1.25 Vpp
    re-align camera+intensifier to electrodes
    Try to trap!?

    2021-03-25 Thursday

    17:04Z: Install all shields, and leave RF on (will switch off after midnight). RF drive voltage: 4.0 Vpp.


    14:16Z Try to load, 1.65 A

    13:41Z Adjust ASE,try to load, 1.65 A (1.66W).

    13:14Z Adjust A-beam, try to load, 1.65 A (1.66 W). No ion.

    12:49Z Adjust A-beam, try to load, 1.65 A (1.66 W). No ion.

    12:10Z Adjust A-beam, try to load, 1.65 A (1.66 W). No ion.

    12:01Z: Try to load, 1.65 A (1.66 W). No ion.

    11:48Z: Try to load, 1.65 A (1.66 W). No ion after 10 min.
    Install intensifier and ethernet camera.

    10:41:30Z: Try to load, 1.65 A (1.66 W). No ion after 10 min.

    10:20Z: Try to load, 1.6 A (1.56 W), RF drive 3.9 Vpp, CH4 monitor 3.02 Vpp. No ion after 10 min.

    Check power: 422 A: 1.1 uW B: 1.5 uW , 461: 144 uW , 1033: 55 uW, 1092: 10 mW

    Check 461 nm dispenser: T = 4.642 kOhm, I = -197.27 mA

    Added 3 dB attenuation. Now 'normal' drive voltage is 3.9 Vpp (CH4 monitor 3.02 Vpp), and high is 8.9 Vpp (CH4 monitor 6.77 Vpp). Voltage can be adjusted now between normal and high without glitching.

    RF generator glitches when moving from high voltage (6.4 Vpp) to 'normal' voltage (3.2 Vpp)

    Logging monitor-voltages from scope failed during the night

    2021-03-24 Wednesday

    Resonator connection, filter board and Cu tube shield had to be removed to put thermal paste between alumina insulator and shapal heat sink.
    Significant drop in monitor voltage after re-assembly was due to resonator helix being off center, towards the trap. After centering (tuner end opened), resonance frequency was <14.414 MHz.
    After tuning, min SWR 1.01 @ 14.424 MHz, R 50.6 Ohm, X -0.24 Ohm. Q visually as before.
    Monitor voltages even higher than before. Note: old monitor very sensitive to helix position, new monitor reflects real trap voltage better?

    2021-03-23 Tuesday

    12:42Z set to 6.4 Vpp, which gives roughly doubled monitor voltages compared to previous night
    12:37Z With 6Vpp drive get monitors of 5.71 Vpp and 2.6 Vpp
    12:33Z RF PA on and connected to helres with 3Vpp drive from siggen CH1

    Test the RF PA. Attenuation from siggen to PA-input: 35 dB
    Previously 70 mVpp (-19 dBm) from siggen gave ca 1 and 2 MHz seculars.
    With 10Vpp (+24dBm) an attenuation of -35 dB gives -11 dBm (ca 178 mVpp) - OK max voltage at PA input?
    siggen 3.1Vpp, 35 dB attenuation, and RF PA seems to give similar monitor-voltages as before 2.93 Vpp and 1.33 Vpp

    2021-03-22 Monday

    Temperature rise with shapal heatsink and new Cu-tube.
    RF 10Vpp from signal-generator, giving ca 3 Vpp on helres monitor. Seculars not measured since last helres/cu-tube changes, probably ca 0.9/1.8 MHz as usual


    Trap efficiency was now estimated from a 3D FEM simulation (KH 2021-02/03), giving:
    eta_in,z = 0.7264
    eta_in_r = 0.7263
    eta_out,z = 0.6978
    eta_out_r = 0.6976
    
    This can be compared to the old 2D cylindrically symmetric FEM result (AW 23.6.2019 and 9.11.2019, TLi notes 13.11.2019):
    eta_in = 0.75
    eta_out = 0.67
    

    2021-03-18 Thursday

    Installed shapal-rod heatsink, new D42mm Cu-tube, new U-shaped filter-board (with MMCX-connector for tickler)
    Black tape on RF-parts for IR imaging

    measured with DMM: V1 = bottom outer electrode = -5.215 V
    V2 = top outer electrode = -4.739 V
    ARTIQ settings are V12cm = -5V, V12diff = 0.239 V (?check if these are the ones last used?).

    2021-03-08 Wednesday

    Restarting clock laser to try new comb rep-rate piezo.
    PDH settings were lost. These seem to work: Freq 25 MHz, amplitude 1.8 dBm, phase 51.4deg, lock level 0.0. (ARC gain notes 2020-11-21)
    Composite-VGA converter for cavity camera is fried. Camera works ok with portable mini-screen.

    New 1-channel Thorlabs piezo drive for Menlo comb rep-rate lock. Seems to work (check beat-note noise and compare to old data?)

    2021-03-08 Monday

    buy more tape for label-maker: Dymo D1 (12mm, 0.5")

    14:45Z Open cavity doors
    Arrange cavity-rack in preparation for moving clock-laser seed there

    2021-03-05 Friday

    shut down lasers
    ion seems lost, storage time ca 11 days
    stop 405nm after 10 min - still no ion
    Re-lock to Rb-line, seems ok
    Try all beams on and 405 nm on, repumper + clearout on
    07:58Z stop B-count program.
    ion went dark at some point, 04:36 local?

    2021-03-04 Thursday

    17:07 start B-count logging (with B-beam)
    Datafile: 2021-03-04-counts


    Abort A-count logging at 17:03 local.

    Results from A beam fluorescence count monitoring:

    Counts and ambient temperature (110 PT100):

    Correlation (XY-plot) of counts and temperature:

    2021-03-03 Wednesday

    Leave A-beam fluorescence monitor over night. start ca 15:47Z
    gate 1000 ms. Should run for 24 hours.
    Datafile: 2021-03-03-counts2. Note 1000ms gate-time is accurate, but there is dead-time between each data point.


    09:30Z people in the lab again, may cause temperature-oscillation?

    Could be related to 422nm polarization issues also?
    C-beam counts, hands-off before t=300

    B-beam counts when heating the fiber. Hands-off a bit before t=200. Towards end camera-image is quite weak also

    A-beam counts when slightly heating the 422laser -> ion-breadboard fiber. A-beam has polarizer, which converts polarization-rotation to power fluctuations(?)

    Seems low, did voltage-changes move ion out of beam-center?
    Fluorescence in 10ms: A 53.76/1.82, B 22.64/1.33, C 61.05/1.95

    Try again: A 12%, B 2.9%, C 2.9%
    At this point uMotion: A 42%, B 49%, C 48% - user ERROR, wrong voltages applied
    Optimal voltages (Common mode, diff): [ 0.23901015 28.08731594 48.19857708], V12cm = -5 V
    ID13189 uMotion scan 2x2x2
    uMotion A 23%, B 29%, C 32%
    Voltages now at -5 0.2347 28 50.051

    ID13185 z with A-beam
    ID13184 z with C-beam
    ID13183 rhi with A-beam - broad resonance
    ID13182 rhi with C-beam
    ID13181 rlo with A-beam
    ID13180 rlo with C-beam

    A-beam seculars
    z lorentzian fit may not be very good?

    C-beam seculars

    B-beam seculars:

    ID13179 rlo scan 935-939 kHz
    ID13178 rhi scan 980-984 kHz
    ID13177 B-beam z-scan with 20dB, fit finally works. peak height ca 40%
    B-beam tickler z-secular with 17 dB gives 60% high peak - and fit fails - maybe more attenuation needed
    lorentzian fit to z-secular with A-beam seems difficult..
    ID13172 try z-secular with A beam again, 27 dB attenuation - but fit fails again.
    ion still stored

    2021-03-02 Tuesday

    No experiments during the night
    Leave B-beam on.

    Summary: Seculars with -5 0.2347 28 50.051. A-beam
    axial/radial ratios ??

    z-scans where lorentzian fit failed:
    Note these are also asymmetric/hysteretic. we sweep the tickler from low to high frequencies


    ID13170 try again with 23 dB DDS attenuation (up from 17 dB) - fit fails for the 3rd time
    ID13169 narrower range 1843-1844.5 kHz - too strong tickler gives blurry image? may not give lorentzian resonance? - fit fails again!??
    ID13168 fit fails :(
    ID13168 z scan 1842-1846 kHz - peak up to 50% with blurry camera image. As predicted A cools z-secular badly, so resonance is narrow
    z search 1840-1860 kHz - found at 1844!
    old z(-6.1)=1876 kHz and z(-6.2)=1879 kHz, so expect z(-5) at -33kHz around 1843?

    ID13166 rhi scan 980-984 kHz - as before, rhi resonance is broad with A-beam
    rhi search 970-990 kHz - found around 982 kHz
    previous rhi was 954 kHz, so expect it around 985 kHz maybe?

    ID13164 rlo scan 935-939 kHz
    930-940 kHz - peak at ca 936.5 kHz - ca +31.5 kHz from last value with V12cm=-6.1V
    search 920-930 kHz - no peak
    search 910-920 kHz - no peak
    tickler search 900-910 kHz (previous rlo was 905 kHz) - no peak found. 0.1V change in V12cm moved r-secular 3 kHz yesterday, so expect 30-33 kHz now?
    Using A-beam for these tickler measurements

    Go back to -5 0.2347 28 50.051 - uMotions A 2%, B 9.9%, C 13%
    At this point: A 46% !!, B 2.3%, C 16% - repeat A 46%
    Optimal voltages (Common mode, diff): [ 0.19201393 27.06730014 49.21492668]
    ID13146 uMotion scan 2x2x2 grid
    at -5 0.2347 28 50.051 get uMotion A 2.8%, B 11.5%, C 12%
    Try scanning V34cm, A-minimum at 28 V
    Now uMotion A 38%, B 27%, C 31%
    Try reducing V12cm to -5 V. Other voltages as before 0.2347 26.7769 50.051


    ID13112 uMotion scan with 5x5x5 grid
    Optimal voltages (Common mode, diff): [ 0.23465246 26.7769039 50.05097994], V12cm=-6.1 V
    At this point A 2.9% in 16.3s, B 2.2% in 13.4s, C 2.6% in 21.9s

    ID13106 uMotion scan 2x2x2
    scan gives Optimal voltages (Common mode, diff): [ 0.18054628 26.89479311 50.30482989]
    At this point get A 3.9% in 20.5s, B 12% in 16s, C 4.4% in 19.8s

    at -6.1 0.2277 26.3891 51.2 get uMotion: A 6.7%, B 22%, C 25%
    Sweep V34 diff: A-min at 51.2 V
    uMotion (100kphot): A 70%, B 71%, C 72%
    Fluorescence (10ms): A 74.27/1.48, B 114.38/1.34, C 66.76/1.98 (lab lights are off)

    after ca. 1 min of 405nm irradiation - ion re-appears!?
    re-lock to Rb-line
    08:54Z ion seems dark
    After a stable period, lab temperature oscillates again?? why?

    2021-03-01 Monday

    16:52 leave lab.
    Leave A beam on. No experiments during the night

    Maybe Rb-lock was OK?
    Now A 97.6/2.13, B 38.85/1.62, C 74.11/1.93
    Try re-locking to Rb-line
    repeat A 93/1.9, B 98/1.47, C 72/1.66
    Fluorescence (10ms): A 86/1.6, B 90.67/1.43, C 64.57/1.61

    What if we try to minimize micromotion at V12cm = -6.1 V? Previosly V34cm change was -1.3 V34cm/V12cm
    ID13058 uMotion scan 2x2x2 grid
    B gives really low fluorescence suddenly? 50 seconds to measure 50 kphotons
    Optimal voltages (Common mode, diff): [ 0.22771198 26.38907221 56.11298177] (note previous V34cm was 26.18V so this is now higher?)
    At this point get 100kphot uMotion A 7%, B 2.1%, C 1% (ca 21-22s measurement time with A and C, but with B 102s)
    Tickler measurements now with A-beam
    ID13064 z-scan, with -6.2V was at 1876.4 kHz (B-beam in use, tickler program sets beam!?)
    A-beam fluorescence also quite weak, laser Rb-lock possibly jumped? (to next Rb-line +60 MHz more detuned?)
    Previous peak-height was just over 30%, now get 60% because of less efficient cooling/damping?
    ID13065 rhi scan 952-956 kHz - previously at 953.9 kHz. Now with A-beam, giving very broad resonance (A cools rhi well?)
    ID13066 rlo scan 903-907 kHz - previously at 905.1 kHz. Note FWHM is almost 3-fold narrower



    ID13053 better z-scan
    z scan 1870-1880 kHz - found just above 1876 kHz rhi scan 951-961 kHz - found at 954 kHz
    rlo shifted up by +3 kHz, around 905 kHz
    ID13047 900-904 kHz rlo scan
    Try V12cm = -6.1 V (no change in other voltages)
    uMotion A 8.6%, B 19%, C 21%
    Tickler measurements with B-beam
    axial/radial ratios 1.967 and 2.073


    axial/radial ratios 1.976 abd 2.083
    with V12cm = -6.2 V , seculars today are:

    ID13045 rlo scan
    ID13044 search for rlo 897-907 kHz - 1% higher than Friday would be 902.8 kHz - found at just over 902 kHz
    ID13043 scan 949-953 kHz
    ID13041 rhi search 940-950 kHz. same scaling would give 946.7 kHz - found peak at 951 kHz
    ID13040 1876-1882 scan.
    ca 0.55% higher than on Friday
    ID13039 search 1870-1880 kHz. 10 kHz/30 pts = 333 Hz step - peak around 1879
    Search for z-secular 1860-1870 kHz

    at this point A 4.0%, B 2.2%, C 1.2%
    Optimal voltages (Common mode, diff): [ 0.23807795 26.17776091 57.07942005], V12cm = -6.2
    ID13032 uMotion search 2x2x2 grid

    ca 57V seems better, uMotion A 13%, B 40%, C 43%
    previous V34diff was 49.8V
    uMotion A 75%, B 74%, C 72%
    Fluorescence A 122.37/1.85, B 119.98/1.83, C 82.64/1.95
    install inner shield camera side, and mid shield

    With old objective, when focused to Z=7mm, gap to vacuum window is ca 5mm
    However image with new objective not so great :( so go back to old objective
    Try new objective: focus is closer to Z=6mm, and new image ca half screen to the right of old objective (as viewed in ASI camera software)
    For old imaging objective XYZ Z=7mm translator X=55, Y= 36<->0
    got C-beam aligned, via electrode diffraction pattern (only inner shield installed)
    remove mid-shield
    aligning C-beam with mid-shield installed seems difficult
    ca 12-12:30Z (14 local) work in the lab (check how temperature responds?)
    Ion stored for ca 1 week now
    07:13:00Z Stop collision monitor.

    2021-02-26 Friday

    16:03Z start collision monitor. B-beam, detuning 208 MHz, fast 3 ms gate. Threshold set as 40 bright 1 dark (artiq sets geometric mean of these?)
    clearout is on
    NOTE: scripts assume 10ms gate time, use last column (machine unit) for analysis
    states
    dark_durations
    bright_durations


    B fluorescence in 10ms: 157/1.14, in 5 ms: 71.97/0.44, in 3 ms 44.27/0.42 - lab lights are off
    Set RF drive to 10 Vpp
    uMotion A 2.8%, B 2.6%

    Summary of secular results (Compare to 2020-12-09)
    Z-frequency 94.95% of Dec value, with voltage-slope 93.09% of Dec value
    rhi 95.55% of Dec value, with voltage-slope 93.37% of Dec value
    rlo 95.05% of Dec value, with voltage-slope 93.90% of Dec value

    Set RF drive to 9.0 VPP.
    z maybe 1681 kHz? found at 1675 kHz
    rhi maybe 836 kHz? found at 852 kHz
    rlo maybe 800 kHz?


    Set RF drive to 9.5 VPP. Expect 95% of 10Vpp seculars?
    z ca 1775 kHz? actual resonance ca 1771 (94.8%)
    rhi maybe 888 kHz? found at 897 kHz
    rlo maybe 843 kHz? found at 847 kHz


    lower radial slightly lower than expected, below 894 kHz
    Try to find r_hi 934-940 kHz
    Behold, artiq autogenerated figure output from the correlation tickler secular sweep program
    (these measured with the usual 10Vpp RF drive, B-beam, fixed 20 dB physical atteunator on dds-output, software attenuation adds to this)


    ID12911 same scan, modify lorentzian fit to match matlab 'lorentzfit' functional form. resonance at 1868.59
    try 1866-1870 kHz - 22% peak at 1868.6 kHz
    ID12909 secular z-sweep, B-beam, 1862-1866 kHz
    RF monitor 3.06 Vpp

    uMotion A 1.8%, B 3.4% (C not aligned to ion...)
    Set -6.2 0.2445 26.9255 49.8
    scan V34diff, A-optimum at 49.8V
    uMotion: A 37%, B 28%
    C-beam alginment seems lost...
    Fluorescence countsin 10ms are: A 97.8/1, B 106.4/1, C 42.2/1.5 (lab lights are off)
    Detuning back to 208 MHz

    11:59 stop collision monitor
    11:58Z ion still stored. fluorescence quite weak

    2021-02-25 Thursday

    lights off, leave lab.
    16:49Z start collision-monitor ID12879, DB 2021-02-25-run1
    B beam, Counts / 10 ms: Bright: 35, dark: 3.
    Detuning 190 MHz. Clearout on
    states
    dark_durations
    bright_durations


    Set detuning to 190 MHz, giving ca 30-40 counts in 10ms
    Use B-beam (may have polarization issues? correlates with lab-temperature?)

    uMotion A 4.8%, B 5%, C 2.8%
    Set voltages -6.2 0.2445 26.9255 51.6
    Scan V34diff, A-minimum at 51.6V
    uMotion A 12%, B 11%, C 11%
    Counts A 141/4, B 62/3, C 103/4

    Try I1 = -8.6056, I2 = -1.0049, I3 = 7.0059, giving abs(B) =
    ID12845 peak search with 300us
    C1+ 18%, C1- 32%
    C1 B estimate = 2.055935940628926
    C1 Linecenter = 75.97028908312774

    NOTE: B-beam fluorescence down to 53.57/3.17, indicating maybe Rb-unlock? (or polarization-effect giving unbalanced C+ and C- peaks?)

    Try I1 = -8.6056, I2 = -3.0049, I3 = 7.0059, giving abs(B) = 2.2646, 2.2526, and uT
    ID12842 peak search, 300us. Should now compute line-center and B after each pair
    C1+ 24%, C1- 64% (!)
    C1 Linecenter = 75.97024614563541
    C1 B estimate = 2.2646399566052304 uT
    C2+ 24%, C2- 14%
    C2 Linecenter = 75.97025830181504
    C2 B estimate = 2.2525933966917435

    ID12840 peak-search with 300us, and new B-estimate 2.119 uT
    C1+ now lower 29%, and C1- higher 53%
    C1 Linecenter = 75.97016610304007 MHz
    C1 B estimate = 2.1102318715100816 uT
    C2+ 45%, C2- 19%

    C1 peaks

    C2 peaks

    C4 peaks

    re-initialize search with C1 line-center and B-estimate
    ID12838 peak search, 300us probe, with B=2uT estimate
    C1 Linecenter = 75.97022862199493 MHz
    C1 B estimate = 2.119082202980859 uT

    C1 peaks

    C2 peaks

    C4 peaks

    Set coil currents I1 = -8.6056, I2 = -2.0049, I3 = 7.0059
    TLi notes on Magnetic Field (2021-02-15)
    For B-beam, using A-matrix and B0 residual field from 2021-01-21. 
    Should give equal C1/C2/C4
    abs(B)  I1      I2      I3
    2       -8.6056 -2.0049 7.0059
    3       -10.672 -2.5016 9.9934
    4       -12.739 -2.9984 12.981
    

    ID12835: Peak search without OD filters, B beam. clock_probe = 300 us
    C1 Linecenter = 75.97019812711028
    C1 B estimate = 1.046407999323715

    C1 peaks

    C2 peaks

    C4 peaks


    ID12831: Rabi flopping on B beam, no OD filters. tau_pi ~ 300 us.


    Problem with comb rep-rate lock
    The closed-loop rep-rate error signal started to show glitches, with about 400-500ms period
    This caused all SDR counter-data to be flagged invalid
    A Thorlabs MDT693 Piezo-driver was tested - and showed no glitches
    Also the cable from the rep-rate lock-box to the piezo driver was replaced!? (suspect looking ends??)
    With the new cable, also the old Piezo driver seems to work? (since about 15:25 local, at least for about 10 minutes)
    Leave it running, to see if problem re-appears
    however the problem re-appears after a 10min warm-up period!!??
    14:08Z Go back to the Thorlabs driver. This gives glitch-free lock and valid SDR-counter data

    after lunch fluorescence seems weak. Re-locking to Rb-lines improves counts 3-fold

    At this point get: A 1.3% in 11s, B 0.9% in 15s, C 0.8% in 15s
    Optimal voltages (Common mode, diff): [ 0.24445048 26.92545019 51.96431002]
    ID12794 uMotion search 4x4x4 grid

    ID12793 repeat - [2.90072747e-02 2.89735223e-04 1.86382763e+00 2.69222756e-04], Lorentzian FWHM = 2 g = 538.4 Hz
    ID12792 scan 1862-1866 - 30% just below 1864 kHz - Lorentzian fit: [2.33235839e-02 3.39371929e-04 1.86379851e+00 3.83917947e-04] (constant, amplitude, f_center, gamma)
    ID12781 search for z-resonance at higher frequencies 1856-1866 kHz
    ID12790 tickler-correlation, z - no resonance


    These can be cropped to show only the ion by:
    convert 2021-02-25_A_15s.jpg -crop 150x150+1050+120 A4.jpg

    Ion images with different beam combinations:
    15s stack of 1s exposures. Old imaging objective, ASI camera, high-gain, no binning (1936x1096 pixels). Pixel size is 2.9 um x 2.9 um. Estimated magnification is 30x
    A
    B
    C
    AC
    BC
    AB
    ABC
    Fluorescence levels (10ms): A 132.1/3.8, B 127.6/4.2, C 75.8/4.0
    Multiple beams: AB 117.7/4.2, BC 154.4/4.3, AC 143.0/4.6, ABC 181.85/4.6
    uMotion A 2%, B 0.9%, C 6.1%
    Set -6.2 0.23 26.8378 52.4
    sweeping V34diff shows A-minimum around 52.5V, finer resolution sweep gives 52.4V
    uMotion A 47%, B 45%, C 49%
    Voltages now: -6.2 0.23 26.8378 55

    Increased AC-setting at 30 Hz gives slightly higher amplitude oscillations, compared to weekend setting of 25 Hz
    Weekend had less heat load in the lab

    07:37 stop collision monitor
    07:36Z ion still stored

    2021-02-24 Wednesday

    ID12763: Start at 14:48:52 UTZ: Collision monitor with clearout on, with B beam. Tag: 2021-02-24-run1
    Counts / 10 ms: Bright: 121, dark: 3.
    usual detuning for operation, 208 MHz (check?)
    states
    dark_durations
    bright_durations


    AC setting now increased to 30 Hz

    Sweeping V34_diff gives new 'optimum':
    uMotion A 4.1%, B 1.2%, C 6.1% at [-6.2, 0.23, 26.8378, 55] (down 1.6V from morning value of 56.6V)

    uMotion A 30% in 12s, B 30%, C 36% at [-6.2, 0.23, 26.8378, 56.6]
    re-lock to Rb-line -> fluorescence restored!
    14:33 ion seems dark/gone?

    ID12745 test2 of secular scan Lorentzian fit: [2.00147896e-02 2.50019692e-04 1.85449329e+00 3.99132412e-04]

    ID12744 test of secular scan with constant+Lorentzian: [2.17395264e-02 2.64289160e-04 1.85434871e+00 4.59689160e-04] (constant, amplitude, f_center, gamma)
    Compare this to Matlab fit to same data points?

    Try the third (lowest) piezo-channel of the driver. This again initially gives ok comb-lock and beat-note data - but after 1-2 minutes glitches appear again!
    rebooting the rep-rate piezo driver seems to remove the glitch - however the problem appears after a few minutes again
    Comb rep-reate lock shows 400ms heart-beat error, can't determine line-center from comb data....
    ID12736 carrier C1 peaks with 600us probe
    +C1 ? see two peaks?
    Now optical carrier and side-bands

    ID12739 z 1.848 - 1.856 MHz, Lorentzian fit: [3.72032165e-04 1.85437750e+00 6.67644320e-04]
    ID12738 r_hi 937-943 kHz,
    ID12737 r_lo 885-892 kHz,
    (these over lunch). more points and 100 kphotons per point

    ID12734 r_lo 885 to 891 kHz, 7 dB - 22% peak at ca 889.6 kHz - Lorentzian fit: [3.85734824e-04 8.89606990e-01 5.82098050e-04]
    ID12733 r_hi 937 to 943 kHz, 7 dB - 16% peak at ca 929.1 kHz - Lorentzian fit: [2.93961269e-04 9.39172793e-01 7.09967596e-04]
    ID12731 z 1849 to 1855 kHz, 17dB - 22% peak at ca 1852.1 kHz - Lorentzian fit: [3.88708773e-04 1.85215658e+00 6.40710257e-04] (curret ARTIQ fit doesn't have constant background term)
    ID12730 r_lo 891 to 897 kHz - no clear peak
    ID12729 secular correlation search 943 to 948 kHz, 7 dB att - no clear peak
    B-beam in use, 50 kphotons/point
    RF monitor ca 3.040 Vpp (slightly lower than y-day?)

    Fluorescence levels (10ms): A 139.2/1.56, B 126.46/1.1, C 110.8 / 1.9
    At -6.2 0.23 26.8378 56.6 get: A 1.4%, B 3.1%, C 8%
    Scanning V34diff shows A-minimum at 56.6 (scanning only V34diff seems to be a quick way to find the optimum from day to day!?)
    Maybe V34diff has settled to a lower value again?
    uMotion 100kphot: A 74% in 9s, B 73% in 9s, 75% in 9s
    Voltages -6.2 0.23 26.8378 63.3441
    08:23Z ion still stored. secular sideband measurements completer over night

    2021-02-23 Tuesday

    15:21Z leave lab

    ID12705 z, search21
    ID12704 r_hi, search20
    ID12702 B-beam r_lo, search19
    ID12701 switch to B-beam
    ID12700 z, 2 kHz AOM span, search18
    ID12699 r_hi, 2 kHz AOM span, search17
    ID12698 r_lo sidebands, 2 kHz AOM span, search16
    ID12697 z sidebands, search15
    ID12696 r_hi sidebands, search14
    ID12695 A-beam, r_lo sidebands, wider 1 kHz AOM span, 16 Hz AOM step, C1+ and C1-, 5 ms probe (no OD filter). DB: search13
    63 points per peak
    C1+ red appears far right in span?

    ID12690 r_lo sidebands

    uMotion A 66%, B 64%, C (very weak)

    ID12676: Photon correlation secular sweep for higher radial: 20 + 7 dB att. Peak at approx. 0.9447 MHz.
    ID12675: Photon correlation secular sweep for lower radial: 20 + 7 dB att. Peak at approx. 0.8942586 MHz.
    ID12674: Photon correlation secular sweep for axial: 20 + 17 dB att. Peak at approx. 1.8524828 MHz.

    ID12673: Secular sweep with 0dB + 22 dB: Lower radial ~ 0.89449 MHz, upper radial ~ 0.94558 MHz
    ID12672: Secular sweep with 0dB + 20 dB: Axial: peak at approx. 1.8528 MHz

    ID12665: Carrier search for C1, tau_pi = 600 us. 2021-02-23-search11-


    ID12659: Carrier search for C1, tau_pi = 600 us. 2021-02-23-search7-

    12657: C1+ red low radial?
    12658: Rabi on previous. Tau_pi maybe 5000 us
    ID12654: Secular search for C1, 2021-02-23-search4-

    Tests with objective, umotion again high! V34diff needs 10 more Volts to get uMotion acceptable (sharp camera image)
    New imaging objevtive did not show a good image... further testing required

    Here uMotion A 0.3%, B 1.1%, C 2.5%
    Back to -6.2 0.2358 26.8378 48.3441

    ID12569 tickler 20+17 dB, 1864 - 1868 kHz, 30 points - no peak..

    At these voltages get 100kphot: A 9.2% in 9s, B 3.4% in 10s, C 5% in 12s
    Optimal voltages (Common mode, diff): [ 0.2358202 26.83783696 48.34408991]
    Test how trap parameters change at V12cm of -6.0V (previous setting was -6.2V 0.2289 26.8187 48.4554)
    ID12562 uMotion scan 4x4x4 (lunch) - USER ERROR: this scan at V12cm=-6.2V as previously....

    Note matlab lorentzian fit is variant '3c' of lorentzfit:
    L3C(X) = P1./((X - P2).^2 + P3) + C
    FWHM is computed as:
    FWHM = 2 * sqrt( P3 )

    fine-resolution correlation measurement of seculars (now with B-beam)
    50 kphotons/point
    ID12555 scan 890-893 kHz, 30 points, 20 + 7 dB attenuation - peak height up to 40% contrast
    ID12556 scan 937-941 kHz, 30 points, 20 + 7 dB attenuation - peak height 14 % (upper radial is harder to excite with tickler - as before)
    ID12557 scan 1.864 - 1.868 MHz, 20 + 17 dB attenuation - peak height 30%
    RF monitor during these 3.045 - 3.048 Vpp

    search for seculars (A beam)
    tickler 0 dB + 15 dB attenuation - no clear fluorescence change. search 0.9 to 1.1 MHz
    0 dB + 10 dB attenuation
    0 dB + 5 dB attenuation. one mode at 939.6 kHz
    0 dB + 2 dB attenuation: lower at 891.4 kHz, upper at 939 kHz
    0 dB + 20 dB for axial, 1.8-2.0 scan, at 1.86644 MHz

    500kphot A 1.8% in 41s, B 1.3% in 49s, C 3.7% in 61s
    100kphot A 1.3% in 9s, B 0.9% in 10s, C 2.9% in 13s
    found Optimal voltages (Common mode, diff): [ 0.22887765 26.81872373 48.45535893]
    ID12530 2x2x2 uMotion search
    Voltages at -6.2 0.2832 27 47
    uMotion 100kphot: A 17.6% in 9s, B 17.3% in 10s, C 32% in 12s
    Fluorescence levels (10 ms): A 160.7 / 2.14, B 133.59 / 1.94, C 98.25 / 2.04
    back to 208 MHz detuning

    08:31Z stop collision monitor

    2021-02-22 Monday


    ID12517: Start at 15:00 UTZ: Collision monitor with clearout on, with all beams, detuning = 190 MHz. Tag: 2021-02-22-run1
    Counts / 10 ms: Bright: 79, dark: 3.
    states
    dark_durations
    bright_durations


    reducing RF-voltage to 4.1Vpp doesn't show two ions - so this is probably a single ion
    These are higher than previously during collision-monitoring/clock-runs (ca 60-70 bright counts?)
    10ms fluorescence levels A 166.5/2.25, B 145.54/1.94, C 109.88/2.54

    all beams ABC show fuzzy image, any single beam sharp image?
    uMotion A 14% in 9s, B 9% in 11s, C 22% in 11s
    14:39:30 signal (in 3 min)
    bias voltages -6.2 0.2832 27 47
    RF back to 10Vpp
    14:36:30Z try load with 1.60A (1.54W)

    changing V12cm from -6.2V to 0V in a single step threw the ions away
    reduce RF to 4.0V increases ion-separation clearly on camera
    14:30:05Z signal, blurry at first, then two separated ions
    14:23Z try to load with 1.65A. detuning 185 MHz, 405nm was up to 7mW - reduce to 3 mW


    lowering RF-voltage shows up to three ions! (saved some images)

    detuning to 208 MHz
    14:16Z signal! fuzzy image
    back to voltages -6.2 0.2832 27 47
    14:11:15Z load again, 1.7A, detuning to 185 MHz

    adjusting RF-voltage down -> maybe two ions in the trap? lost both at RF-glitch
    bright ion seems to jump between two positions?? (when switching beams) do we have two ions trapped?
    -6.2 0.2831 27 42.4 gives A 15%, B 2.2%
    -6.2 0.2832 27 Scan[41 43] B
    -6.2 0.2832 27 Scan[38, 43] gives A-optimum at 42V

    A 65%, B 68%, C 70%
    Optimal voltages (Common mode, diff): [ 0.2778 26.13 47.58]
    ID12458 2x2x2 uMotion scan
    [0.25, 0.32] [26, 27.5] [46.5 48.5](too large?)

    uMotion A 70% in 8s, B 64% in 11s, C 70% in 11s
    Voltages -6.2 0.2832 27 47
    Lowering trap RF to 5.7Vpp gives stable high fluorescence, 1200 couns / 50ms all beams
    try different V34diff voltages, but no clear effect
    seem to have something trapped, but fluorescence comes and goes
    13:39:35 signal, but fuzzy image of ion? gone quickly?
    13:35:30 try to load, 1.7A, detuning to 185 MHz

    13:32:15 signal! but lost quickly?
    13:26:15 load with dispenser 1.7A. test loading with 208 MHz detuning?!

    back to volgaes ion lost during uMotion search?
    ID12410 seems to have finished ok?
    fit fails because V34diff range was [46.5, 46.5] (no scan at all)
    not so reasonable Optimal voltages (Common mode, diff): [ 1.80241039e-15 9.93688436e-17 -2.15053763e-02]
    ID12410: NPL search 2x2x2.

    Optimize beams and 422 waveplate.
    0.2832 27 47 gives A 14%, B 9%, C 21%
    0.2832 28.1299 47 gives A 28%, B 26%, C 40%
    0.2832 28.1299 46 47 gives A 28%, B 26%, C 40%
    ID12380 scan V34diff vs A-contrast. Minimum at ca 47V
    uMotion 100kphot A 76% in 9.9s, B 70% in 16s, 71% in 15s
    ASE also required re-aligning, got >2-fold more fluorescence (all beams) by aligning ASE
    uMotion 100kphot: A 81% in 22s, B 76% in 27s, C 75% in 29s

    B > 200 counts in 50ms, C was already quite well aligned, up to 280 / 50 ms afer adjustment
    re-aligning beams is required, since V34dif changed som much? 3-fold more A-fluorescece ca 300counts in 50ms
    check Rb-lock, locking to the right line gives better fluorescence
    422 power only 5uW on breadboard before A-fiber (this usually 10-15 uW), adjustment of 422 breadboard fiber-coupling gives ca 9uW
    lowering V34diff to 53V gives the usual diffraction-pattern image of ion
    detuning to 208MHz
    12:05:50 some weak signal? (ca 3 min loading time - good!?) 12:02:45 load with dispenser 1.7A

    405nm adjusted to 3.7 mW
    422nm 185 MHz detuning

    461nm dispenser test, Ractual 4.642 kOhm, -197.32 mA, test against 5A dispenser ca 2 min to see fluorescence.
    RF monitor 3.045 Vpp
    Voltages to last known values: -6.2, 0.2832, 28.1299, 85.6211

    camera focus back to last value
    11:52Z install mid-shield, and imaging objective
    08:20Z turn on stuff to prepare for trapping again

    Menlo oscillator was on continuously.
    A/C VFD was at 25 Hz
    With most stuff turned off on Friday, see strange oscillations in lab temperature

    2021-02-19 Friday

    Leave the RF siggen on/off program on
    Clamp down helres trap-side
    Trim 3mm off the helres-mount clamp

    2021-02-18 Thursday

    Before final fine-tuning, vswr with newer rigexpert analyzer:

    Helical resonator min SWR was 1.12 @ 14.422, which probably explains lower seculars (qz). HR moved a few mm to the left and towards the trap and rotated ~0.5…1 deg CW seen from tuner end to center helix and to align shield adapter with trap Cu pipe.
    Also 0.1 mm soft Cu shield replaced by stiffer 0.2mm Cu. This increased Vmon (helix moved closer to antenna) and required tuning and impedance matching.
    After: min SWR = 1.01 @ 14.424 MHz and Vmon = 2.995 Vpp (averaging 32).
    R = 50.2 Ohm, X = 0.15 Ohm, Q ~ 800 as before.

    NOTE: adjusted A/C to 25 Hz - to possibly reduce temperature swings?

    detuning to 185 MHz, 461nm and 405nm lasers on - but no ion?
    08:08Z Ion seems lost? Rb-lock was unlocked (large temperature-swings in lab!?), re-lock, but this doesn't restore fluorescence

    NPL uMotion program had aborted, due to ARTIQ error? (because of Rb-unlock, or ion loss?)

    2021-02-17 Wednesday

    Leave 8x8x8 uMotion scan over night
    ID12326 ranges [0.1 0.4] [26 30] [80 90]

    16:09Z uMotion 100kphot: A 34% in 8.7s, B 31% in 14s, C 24% in 9.7s
    The required V34diff might be drifting down during the day?

    TLi computed from these: qz=0.3643, a=(-0.00038, 0.0016, -0.0012)

    These measured with 10Vpp drive, ca 2.70 Vpp helres monitor, and fixed -6.2 V12cm
    Reults to be plotted + Lorentz-fit: axial ID12318, r-lo ID12320, r-hi ID12322

    upper radial with photon correlation. B-beam
    ID12321 978 to 982 kHz, 20+7 dB att, 10 kphotons
    ID12322 with 50 kphotons/point

    lower radial mode with photon correlation. B-beam
    ID12319 923 to 927 kHz, 20+10 dB att, 10 kphotons
    ID12320 20+7 dB att, 50 kphotons

    fine-resolution secular scan using photon correlation. B-beam. ( 2020-12-09 results use 13 dB for radials and 21-23 dB for axial)
    search areas at 925, 980, and 1892.5 kHz.
    use the physical 20dB attenuator
    ID12312 1.89 to 1.895 MHz in 30 steps, B-beam, 10 kphotons per point, 20+22 dB attenuation
    ID12313 20+26 dB attenuation, 20 kphotons, - too much attenuation, peak height low
    ID12314 20+23 dB attenuation - 0.2 peak height
    ID12316 20+20 dB att, 10 nbins,
    ID12317 20+17 dB att, 0.3 peak height ok? but peak looks a bit noisy?
    ID12318 20+17 dB att, 50 kphotons/point (ca 6-7 s/point)

    Coarse search for secular frequency.
    20dB physical attenuator on tickler-DDS. 0 dB software attenuation
    resonance at ca 1.892 - 1.894 MHz ca 96% of 2020-12-09 result (lower than before, 2020-12-09 result with 10Vpp drive is 1.968 MHz)
    Try 6 dB sw attenuation, resonance at 1.892-1.893 MHz. (2020-12-09 results have radial modes at 985 and 940 kHz, or 50% and 47.8% of axial)
    This would predict current radial modes at 947 and 904 kHz
    tickle 0.85 - 1.0 MHz, 6dB, 400 points, but no clear signal (50ms PMT gate).
    Try 0dB , also try 100ms gate, remove 20dB physical attenuator and use 15 dB sw attenuation
    with 10dB sw-attenuation, one mode at 925 kHz?
    try 5dB attenuation, try 0 dB attenuation -> this reveals upper mode at 980 kHz!

    B fluorescence 75.7 / 3.96 in 10ms (mid-shield installed, back lab lights on)
    B intensity still changes with polarization! Adjust waveplate again.

    uMotions: A: 37%, B: 33%, C: 31%

    Install mid-shield.

    ID12264: Peak search, B-beam, no OD-filters. 2021-02-17-search6-.
    C1 linecenter: 75 971 017.0 Hz
    C1 B estimate: 1.2443346 uT

    C1 peaks

    C2 peaks

    C4 peaks


    ID12258: B-beam Rabi flopping, tau_pi = 0.6 ms:


    Adjusting waveplate on 422 breadboard reduces fluorescence change to negligible levels:


    Heating B or C fibers by hand seems to give >50% fluorescence change

    ID12245 peak search A-beam 350us, no OD filters
    C1+ ca 75 974 7, height ca 40
    C1- 75 967 5, height 60

    Found C1+, rabiflop shows maube 350us tau_pi with A-beam?
    seems Zeeman AOM diffraction-order was wrong. Switching to right one gives quantum-jumps as expected.
    ID12238 spectrum-scan with 1ms probe, no OD-filter, A-beam
    beat-note and excel-sheet suggest line-center around 75.971 150 MHz
    Coil currents (-8, 0, 0) mA - same as previously?
    A-beam fluroescence 97.4 / 2.4 bright/dark in 10ms
    Try to search for the clock-transition?

    At this point get: A 3.8% in 13s, B 2.9% in 10s, C 2.4% in 19s
    Optimal voltages (Common mode, diff): [ 0.2832 28.1299 85.6211]
    ID12223 Try 2x2x2 uMotion scan?

    V0 - 0.33VA = 0.2610 27.8855 87.2692, here get A 27.5 %, B 29.8%, C 36.3% - ??
    maybe -sign is right?
    V0 + 0.33 VA = 0.2612 28.1501 87.5210 here get A 39%, B 29.6%, C 35.5 %
    Try moving along voltage-vectors to zero uMotion?
    uMotion 100kphot: A 33.8% in 15s, B 29.4% in 12s, C 36.7% in 13s

    detuning constant at 208 MHz
    Rb-signal has a lot of DC-offset, so this may give some extra detuning?
    Re-lock to Rb-line, this gives ca 5-fold more fluorescence and brighter camera image
    uMotion 100kphot A 9.7% in 91 seconds!?, B 9.0% in 95s, 11% in 85s
    Fluorescence level seems low? camera image weak..

    Comb f0 unlocked due to low SNR. re-locked
    (lunch)

    ID12186: 4x4x4 NPL search around previous value.
    Result: [ 0.26114758 28.01782364 87.39507497]
    uMotions at optimum: A: 3.6%, B: 5.5% C: 9.2%
    Check that the columns of inv(Cfit) are voltage-vectors that change only one contrast!?
    % Micromotion contrast model
    % S = C (V - V0)
    % S has [A, B, C] contrasts
    % V and V0 have voltages [V12dif, V34cm, V34dif]
    % V12cm is constant -6.2 V in recent experiments
    V0 =
        0.2611
       28.0178
       87.3951
    Cfit =
       -0.9303   -2.0943   -0.4214
       -0.2217    0.1219   -0.1284
       -0.1447   -0.1542    0.1619
    inv(-1*Cfit) =  % sign is arbitrary..
       -0.0004    2.9735    2.3561
        0.4009   -1.5572   -0.1910
        0.3815    1.1746   -4.2519
    Example:
    Cfit * [-0.0004 0.4009 0.3815]' = [-1 0 0]'
    moving along first column-vector of inv(Cfit) changes A by 100% and leaves B and C unchanged.
    Cfit * [2.9735 -1.5572 1.1746]' = [0 -1 0]'
    Cfit * [2.3561 -0.1910 -4.2519]' = [0 0 -1]'
    

    Manually: 29%, 30%, 31% at [0.25, 29.4, 87]
    Manually: A: 53%, B: 1%, C: 21% at [0.3306, 27, 86]

    2021-02-16 Tuesday

    ID12094 leave uMotion scan over night
    8x8x8 grid. voltage ranges [0.5, 1.5], [20, 30], [95, 105]
    Result: (Common mode, diff): [ 3.99139299 12.57404019 85.31937744]

    1 scanned 25.8, 98: A-optimal (A 1.9%, B 74%, C 40%)
    1 25 99 seems roughly A-optimal (A 7% B 65% C 8.4%)
    go back to the AC-optimal 1.0 25 100: A 9.7%, B blur 62%, C 7.1% - so AC-optimal point is roughly repeatable

    but scanning V34diff or V34cm around here gives more or less the same flat A-contrast values
    at 0.3 27 48 get A 21% B 29% C 26%
    at 1.0 27 50 get A 22%, B 32%, C 26% - images look blurry, but this may be due to focusing of camera? or beam-alignment needed?
    These voltages are far from the 2021-02-03 values [0.28, 27.0, 48.0]...

    B-blur better at 1.0 32 102.5 (A blur 57%, B 77%, C 58%)
    C-nubunyn at 1.0 28, scanned 102.5 (A 78%, B blur 64%, C 4.1%)
    B-blur is sideways, so change C34cm? at 1.0 28 100.2 get (A 71% ,B 78%, C 46%) all images sharp
    C-minimum at 1.0, 25, scanned 100.2 (A 1.9% !, B blur 63%, C 0.3%) first found point that is AC-optimal
    B-blur better at 1.0, 25, 98.2 (A 45%, B 77%, C 45%)
    C-minimum at 1.0, 22, scanned 98.2 (A 75%, B 63% blurry sideways, C 5.2%)
    A-blur reduced at 1.0, 22, 96.2 (A 75%, B 75%, C 36%)
    C-contrast minimum at 0.6, 22, scanned 96.2 (A 59% blurry, B 66%, C 1.3 %)
    C-contrast minimum at 0.6, 23, scanned 97.4 (A 77%, B 68%, C 3.7%)
    A-blur reduced at 0.6, 23, 96.44 (A 75%, B 61%, C 20% )
    C-contrast minimum at 0.3543, 23 and scanned 96.44 (A 61,6% blurry vertical image!, B 19.5%, C 4.6%) C-contrast minimum at 0.3543, 24 and scanned 97.33 (A 75,6%, 19.4 %)
    C-contrast minimum at 0.3543, 23.4963 and scanned 96.88

    Manually: A: 75%, B: 27%, C: 13% at [0.3306, 23.455, 97.98685]

    Further searching gives 0.3, 29, 93 V. This gives reasonably sharp ion image with clear diffraction rings
    Here uMotion A 75%, B 77%, C 76% - correlation histograms roughly sine-shaped

    Manually found better voltage combination for A and B (harmonic, but high contrast): [0.9, 28, 91]

    detuning changes fluorescence very little - indicating lots of uMotion?
    detuning back to 208 MHz (from loading value of 185 MHz)
    similar blurred image as before - showing blurred diffraction-rings typical of the imaging objective
    blocking some beam briefly gives a cooled ion
    12:48:05Z signal (axial mode? not cooled?)
    12:43:15Z load with 1.7A dispenser

    12:40 throw away old ion(s)
    lots of uMotion? maybe have 2 ions?

    ID11722: Leave NPL umotion search over lunch.

    Ion image blurry, lot of micromotion?
    Blink test shows only 2 levels, single ion?
    Readjust PMT splitter.

    try to align all beams, so we get bright fluorescence
    422 detuning to 208 MHz
    07:44 signal! quite far right from where camera center was aligned?
    07:43 load with 1.7A dispenser
    align A-beam with help of output diffraction pattern
    remove PMT-splitter, align camera to electrodes

    07:31Z dispenser off, no clear signal
    07:18Z try to load, 1.7A dispenser.
    Adjust 461 current.

    06:55Z try to load, 1.7A dispenser. No ion after 15 min.
    Adjust A-beam position

    2021-02-15 Monday

    Cavity drift with minus-K floating now seems better? This day Lunch 12-13 and team-development 13-15 - chronograf figure with local time on X-axis?


    need to remove PMT-splitter, to get better image with intensifier?
    15:04Z no signal..
    intensifier gain 3.77V, PCO camera exposure 222ms
    14:53Z try to load, 1.7A dispenser.
    install intensifier + PCO-camera

    need to install intensifier? to align A-beam?
    14:19Z no signal in 14 min.
    14:05Z try to load, dispenser 1.7A (1.75W.. rises to 1.815W after 10min)

    ASI camera USB-cable pointing vertically down gives horizontal inner-electrode images
    with modified imaging-objective (0.5" slimmer lens-tube, to fit outer shield cimney), focus position is ca 1mm back from previous setting
    then move in Z 425um up. (upper electrode appears ca 850um up from lower)
    focus on lower electrode, then move focus inside 750um/2.
    removed also mid and inner shield (camera side) in order to find electrode images
    remove outer shield, to align camera to electrodes

    no signal in 11mins... check camera alignment to electrodes?
    09:00Z try to load, with 1.7A (1.74 W)
    RF monitor 2.74 Vpp

    Beam powers:
    A 461nm, 250 uW (fiber was blocking IR-beam, and temp/current adjusted after moving driver-boxes)
    422nm A 0.9uW, B and C 1.0 uW, detuning 185 MHz
    405nm A adjusted to 3 mW

    Check 461 driver settings: T = 4.642 kOhm , I = -195.23 mA (dispenser 5A and ca 2min as usual)

    Adjusted 422 input polarization.

    Temperature rise with outer top shield installed


    2021-02-08 Monday

    14:45Z close cavity-doors after adding 3rd minus-K clamp and aligning laser to cavity

    SRS shutter test:
    Quiet: 10:06 - 10:19 (STDEV: CH1: 26.4 Hz, CH2: 26.4 Hz), SRS shutter (1 ms probe): 10:21 - 10:34 (STDEV: CH1: 27.4 Hz, CH2: 27.4 Hz).

    2021-02-04 Friday

    14:47Z close cavity doors
    Clamp down minus-K so it doesn't move

    turn off cooling-laser, ASE.
    no ion.. Storage time from Monday 25.1 to Friday 4.2, ca 11 days?
    try large 422-detuning, and ramping back to 208 MHz
    try 461nm laser
    try 405nm laser
    try to re-lock to Rb-line, but no ion..

    Shutter noise tests:
    Nothing happens: 15:14 - 15:19
    Dummy clock run with 1000 ms probe: 15:20 - 15:25

    Ion seems to been lost.

    ID11687 NPL uMotion search, 100kPhotons. Result: [ 0.24382725 26.82098208 48.19374483]


    ID11686 Peak search with OD2+OD3 based on C4 linecenter and B estimates from clock run. Probe = 190 ms, 2021-02-05-search-1
    0.4 Hz sampling, 13 Hz range. 2 Hz expected AOM FWHM.


    C2 peaks

    C4 peaks


    Switch to OD2+OD3. Try to operate experiment remotely.
    10:16Z Abort clock-run

    2021-02-04 Thursday

    ID11684: Clockrun with OD1+OD3, 60ms probe, based on C1 estimates of linecenter and B field. 2021-02-04_clockrun1
    New compared to last clock-run: Continuous drift-compensation with ARTIQ-DDS, adjusted in real-time to 2h drift-estimate from comb data. 2nd-order servo in peak-lock algorithm.

    ID11669: OD1+OD3 peak search with probe = 60 ms. 2021-02-04-search18-. Initialized based on C1 linecenter and B from previous measurement.
    C1+ 75 965 302.4 Hz, FWHM 5.6 Hz, h33
    C1- 75 959 600.5 Hz, FWHM 7.0 Hz, h43
    C1center = 75 962 451.4 Hz, B = 1.016029 uT
    C1 peaks


    ID11678: Rabi flopping on C1+ with OD1+OD3. ~60ms still ok.


    ID11669: OD1+OD3 peak search with probe = 60 ms. 2021-02-04-search17-. Initialized based on C1 linecenter and B from previous measurement.
    C1+ Center visually 75.965301
    .

    ID11673: Rabi flopping on C1+ with OD1+OD3. Excitation only ~16% at max.

    ID11671: OD2+OD3 peak search with probe = 190 ms. 2021-02-04-search15-. Initialized based on C4 linecenter and B from previous measurement.
    Sampling 0.4 Hz, range 12 Hz. Aborted during C2+.

    C1 peaks

    C2 peaks

    ID11669: OD1+OD3 peak search with probe = 60 ms. 2021-02-04-search14-. Initialized based on C1 linecenter and B from previous measurement.
    C1+ 75 965 311.5 Hz, FWHM 7.4 Hz, h30
    C1- 75 959 609.5 Hz, FWHM 6.8 Hz, h44
    C2+ 75 968 143.3 Hz, FWHM 6.9 Hz, h38
    C2- 75 956 774.3 Hz, FWHM 7.9 Hz, h26
    C4+ 75 976 676.1 Hz, FWHM 8.4 Hz, h39
    C4- 75 948 237.5 Hz, Unreliable fit, noise in the data.

    C1center = 75 962 4605.2 Hz, B = 1.016047 uT
    C1 peaks

    C2 peaks

    C4 peaks



    ID11667: OD1+OD3 peak search with probe = 60 ms. 2021-02-04-search13-. Initialized based on C1 linecenter and B from previous measurement.
    C1+ fit failed. Visually 75.965312
    Linecenter 75.9624608364173 MHz.
    B 1.016095

    ID11664: OD1 peak search with probe = 1.9 ms. 2021-02-04-search11-
    C1center = 75.96246910765049 uT
    B = 1.017961 uT
    Abort after C1 pair.

    ID11648: OD1 peak search with probe = 1.9 ms. 2021-02-04-search2-

    Drift comp seems stabilized.

    C1 peaks

    C2 peaks

    C4 peaks


    ID11643, ID11646: NPL searches with wider range. Trying to debug contrast fitting errors.

    uMotions: A: 4% in 12s, B: 1% in 10s, C: 8% in 17s

    ID11637: Peak search with OD1. Probe 1.9 ms. 2021-02-04-search1-
    +C1 75 965 329.2 Hz, FWHM 240 Hz, h31
    -C1 75 959 617.7 Hz, FWHM 212 Hz, h56

    C1center = 75 962 473.5 Hz, B = 1.0177363 uT

    B-field estimate from the day before (1.018202) seems good enough to predict transition frequencies for OD1.

    C1 peaks

    C2 peaks

    C4 peaks


    Adjust comb f0 lock. Restart drift-comp.

    07:02Z Abort remaining runs. Drift-comp stopped working 2020-02-03 14:20Z.

    2021-02-03 Wednesday

    leave OD5, 190ms scan over night... may or may not find something?
    clock-laser was unlocked for some time at the start of C1+ scan. from pt 33/1000 onward it should be ok
    C1 peaks

    C2 peaks

    C4 peaks


    Go to OD3 and 20ms, to find lost peaks..
    4 Hz step, 200 Hz span, 51 pts/peak
    C1 fit fails... left in the 200Hz span - can we have this much drift?
    C1 peaks

    C2 peaks

    C4 peaks

    scan around 75 965 498, to find C1+?
    find only low-height peak, try rabiflop there? get only low counts, so we are off-resoncance?

    ID11627 Try instead peak-search with OD4 and 60ms probe
    C1+ far left in span, maybe 75 965 498 ? C1- also left? 75 959 780 or lower? (B seems to have decreased? moves C1+ left and C1- right?)
    190ms scan is possible only with well-compensated cavity-drift!
    ID11619 repeat OD5, 190ms peak-search (search9)

    Reset estimates based on C4 linecenter and B-field from OD1+OD3 peak search.

    Estimate tau_pi = 190 ms with OD2+OD3

    ID11618 Peak search with OD2+OD3, 190 ms probe 2021-02-03-search8-.
    Sampling 0.4 Hz, range 12 Hz.

    +C1 75 965 513.20 Hz, FWHM 2.2 Hz, h26
    -C1 75 959 798.60 Hz, FWHM 2.1 Hz, h41
    C2 fit failed. Visually: 75 968 350.4 h37

    C1center = 75 962 655.90 Hz, B = 1.018282389 uT

    C1 peaks

    C2 peaks

    C4 peaks


    ID11613: Rabi-flopping on C1+ OD1+OD3. tau_pi ~60 ms seems ok.


    Reset B-bield and linecenter estimates based on previous C1 peak search.
    ID11612 Peak search with OD1+OD3, 60 ms probe 2021-02-03-search6-.
    +C1 75 965 513.9 Hz, FWHM 6.5 Hz, h32
    -C1 75 959 799.2 Hz, FWHM 7.1 Hz, h41
    +C2 75 968 352.7 Hz, FWHM 7.0 Hz, h34
    -C2 75 956 959.5 Hz, FWHM 7.4 Hz, h29
    +C4 75 976 905.5 Hz, FWHM 7.2 Hz, h36
    -C4 75 948 404.0 Hz, FWHM 7.5 Hz, h
    C1center = 75 962 656.6 Hz, B = 1.0182969 uT
    C4center = 75 962 654.8 Hz, B = 1.0182016 uT
    C1 peaks

    C2 peaks

    C4 peaks


    ID11610 Peak search with OD1+OD3, 60 ms probe 2021-02-03-search5-. Aborted after C1 pair.
    +C1 75 965 513.6 Hz, FWHM 7.5 Hz, h33
    -C1 75 959 798.8 Hz, FWHM 7.0 Hz, h38
    C1center 75 962 656.2 Hz, B1 = 1.0183112 uT

    ID11607 Peak search with OD1, 1.9 ms probe 2021-02-03-search3-.
    +C1 75 965 505.5 Hz, FWHM 196 Hz, h42
    -C1 75 965 505.5 Hz, FWHM 228 Hz, h49
    C1center 75 962 650.3 Hz, B1 = 1.0175206 uT

    uMotions at new 'optimum': A: 2.5% in 11s, B: 1% in 10s, C: 7.7% in 15s
    ID11959: NPL search again. Optimum out of bounds again. Selected manually: [0.28, 27.0, 48.0]
    ID11594: NPL uMotion search around new 'optimum'.
    uMotions at new 'optimum': A: 6.5% in 10s, B: 7.2% in 10s, C: 3.1% in 14s
    ID11589: NPL uMotion search. Optimum out of bounds. Manually selecting: [0.29, 26.6667, 48]
    uMotions: A: 7.4% in 11s, B: 24.5% in 10s, C:22.0% in 15s

    06:50Z Stop clock run.

    2021-02-02 Tuesday

    ID11583 try to start OD4, 60ms clock-run
    Cycle duration is ca 133 seconds. This should roughly match 6 peaks * 2 R/B sides * 100 shots * (100ms probe + 40 ms dead) = 120 s (check what our actual programmed dead-time is?)
    For clock-run results see ion.
    ID11578 peak-search, OD4, 60 ms (Fourier limit 6.7 Hz AOM)
    1.3 Hz step, 30 Hz span, 23 pts/peak
    +C1 75 965 645.2 Hz, FWHM 7.4 Hz, h29
    -C1 75 959 936.7 Hz, FWHM 7.1 Hz, h39 (right in span)
    C1center 75 962 790.9 Hz, B1 = 1.017194 uT
    initialize clock-run with these values. (initializing with a C4 peak search might be better? more accurate abs(B) value)

    ID11577 peak-search, OD3, 19 ms
    2.5 Hz step, 50 Hz span, 21pts/peak
    50/100Hz sidebands not resoled? this is not a good probe length?

    ID11576 peak-search, OD1, 1.9 ms
    now better peak-height with correct tau_pi +C1 75 965 647, FWHM 202 Hz AOM, h38
    -C1 75 959 931, 208Hz, h49
    C1center 75 962 789, B1 = 1.0185 uT
    +C2 75 968 488, 227 Hz, h44
    -C2 75 957 102, 229 Hz, h34
    +C4 75 977 030, 221 Hz, h47
    -C4 75 948 563, 229 Hz, h27

    OD1 rabiflop on center gives 1.9ms tau_pi.
    674nm power in A-beam seems reduced? re-align?
    ID11569 peak search, OD1 and 3ms - but get low peak-height for +C1, so probe length is wrong?
    ID11568 rabiflop on +C1 (75 965 644), with OD3,
    ID11567 peak search, A-beam, OD3, 30ms probe
    2.5 Hz step, 100Hz span, 41pts/peak

    ion dark moment? abort
    ID11563 peak search, A-beam, OD3, 30ms probe

    uMotion at old voltages 100kphot: A 7.3%, B 19%, C 13%

    ID11557: OD1 peak search with adaptive Artiq drift comp. 3 ms probe, 2021-02-02-search4-

    ID11338.png
    uMotions: A: 6% in 13s, B: 22% in 10s, C: 18% in 15s

    ion still stored (now 8 days?)

    2021-02-01 Monday


    Leave 100ms peak-search over night. 0.8 Hz step, 200Hz span, 251 pts/peak
    A-beam, 18 dB Zatt., OD4, 100ms probe
    C1 peaks

    C2 peaks

    C4 peaks


    Quite difficult to get a clock-run to lock to the peaks with so much uncompensated cavity-drift...
    100ms peak search again. 0.8Hz step 20Hz span
    f_beat_center = 75 962 869
    +C1 far left in span ?

    Try clockrun initialized with (centers, FWHMs AOM, gains) [75.96589229 75.95992971 75.96885552 75.95696648 75.97778133 75.94804067] [4.e-06 4.e-06 4.e-06 4.e-06 4.e-06 4.e-06] [0.5 0.5 0.5 0.5 0.5 0.5]
    100ms probe, 6 peaks gives cycle time of ca 181 s
    somehow we again have +10 mHz/s of uncompensated cavity drift. This gives 1.8 Hz @ 674nm drift during the 180s servo-cycle. For a 100ms probe, 8 Hz Fourier limit, this is 1/4th of the FWHM!

    ID11534 peak search with 100ms
    +C1 far left in span? 75 965 892 (or even lower?)
    -C1 right in span 75 959 929
    C1 center 75 962 911 and B1 1.06247 uT
    predicted line-center from beat-note 75 962 884

    Try clock-run with 100ms probe, but servos seem not to find the peaks..

    Based on OD3 results, switch to OD4 and 100ms probe.
    ID11528 peak search A-beam, 18 dB Zatt, OD4, 100ms probe, 0.8 Hz step, 20 Hz span, 25 pts/peak
    +C1 seems far left in the span? 75 965 901
    -C1 more centered in span (why?),and higher, 75 959 922
    This gives
    C1center 75 962 912 and B1 1.06532 uT

    With OD3 and 30ms, found C1 peaks:
    +C1 at 75 965 910
    -C1 at 75 959 920
    This gives
    C1center 75 962 915 and B1 1.0673 uT
    ID11525 A-beam peak search. 18 dB AOM att, OD3-filter, 30ms probe
    2 Hz step, 200 Hz span. 101 pts/peak
    C1 peaks

    C2 peaks

    C4 peaks


    C beam 106/3 counts in 10ms
    B beam 94/3 counts in 10ms
    A beam fluorescence 173/4 bright/dark counts in 10ms
    422 power A 1.5uW, B 1.9uW, C 1.5 uW
    13:05 after many hours of working with the 422-laser (no cooling), ion re-appeared!

    repeat measurement of the cavity power-shift (Friday result was 190Hz/V @ 1348 nm)
    With reduced FALC-gain, photodiode signal is slightly more than before.

    Cavity drift-compensation should also match true cavity drift better than before
    13:45Z set-point 0.8 V, 5min beat-note 23 164 432.26 (stdev 30.39 Hz)
    13:52 set-point 0.7V (BOA 430 mA), 5 min beat-note 23 164 433.33 (valid data only)
    14:01 set-point 0.8V (BOA 480 mA), 5 min beat-note 23 164 434.19 (31.19 Hz) - not much evidence of power-shift?
    14:24 set-point 0.9V (BOA 520 mA), 5 min beat-note 23 164 434.41 (32.57 Hz)
    14:56 set-point 1.0V (BOA 580 mA), 5 min beat-note (this hits the BOA current limit of 600 mA)

    again dark image after power-measurements? takes a long time to cool and get an image?
    422nm 0.6uW in A-beam
    15-20 uW before ABC-beamsplitter
    A-beam fluorescence 32/12 bright/dark in 10ms gate. B and C weaker
    09:47Z ion re-appeared!? maybe weaker image than before, lower 422nm power after fiber-adjustments
    09:40Z seems we lost the ion?, while adjusting polarization (no cooling-laser)
    422nm Polarization extinction ratio improved from 7 dB to 10 dB, on ion-breadboard after 422-laser-fiber, before ABC-beamsplitter

    Note on cavity doors closed (Friday):
    Coarse heater-power dropped from 90W to 30W - as before. time-constant is 'fast', below one day
    Inner shield free-running temperature sensor seems to settle at to a temperature +4...6 mK than before. Very slow settling, maybe a week?


    09:24Z adjust comb f0-beat SNR to >25 dB
    Updated drift-comp to current observed value of 12.6 mHz/s
    Reducing FALC-gain (trimpot) gives lower servo-bump sidebands and more power transmitted through cavity, now 0.85 V

    'Heating' fiber by hand, both laser -> breadboard and breadboard -> ion.
    ID11492: A-beam, samples: 0-200: laser-> breadboard, samples: 200-350 breadboard->ion, 350 ->: laser to breadboard. Changes mostly in first case.
    ID11490: C-beam, samples: 0-200: laser-> breadboard, samples: 200-350 breadboard->ion, 350 ->: laser to breadboard. Changes in both cases.
    ID11488: B-beam, samples: 0-200: laser-> breadboard, samples: 200-350 breadboard->ion, 350 ->: laser to breadboard. Changes in both cases.
    ID11486: A-beam, PMT count changes with heating laser -> breadboard, but not with breadboard -> laser.
    Check 422 beam polarizations.
    Cancelled long experiment from ID11481 forwards.

    2021-01-31 Sunday


    Leave B-beam, OD5, 60 ms peak search over night
    Maybe 50/100 Hz sidebands are smaller, when t_probe is a multiple of 10 or 20 ms?
    ID11477 (narrow span), 11478(search5), 11479(6), 11480(7) (aborted), 11481(8) (cancelled), 11482(9) (cancelled), 11483 (10)(cancelled)
    C1 peaks

    C2 peaks

    C4 peaks


    Try B-beam instead?
    rabiflop with OD3 suggest maybe 10ms tau_pi?
    peak-search with 10ms and OD3
    OD2 tau_pi looks like 3ms
    OD3 maybe 10 ms
    OD4 no clear oscillation?
    OD5 rise tp 60ms, then flatish

    recalculate peak-positions based on beat-note and latest B1-estimate.
    try peak-search with OD1 and 100ms probe
    +C1 at 75 965 640. try rabiflop there.
    no clear signal, try OD3 and 30ms instead.
    ID11467 peak-search with OD3 and 30ms
    rabiflop on 75 965 669, with OD3. no higher than 25 counts?

    abort long run.

    2021-01-30 kiky-Saturday?

    ID11461 OD3+OD2 and 290ms peak search, long, over night. 290ms is a guess of tau_pi, might be too long?
    18dB Zeeman AOM atteunation.
    0.3 Hz step, 150 Hz span, 501 pts/peak
    duration ca 30s per point = 4 hours per zeeman peak! done in ca 24 hours..
    Fourier limit 2.8Hz optical, 1.4 Hz AOM. ca 5 steps of 0.3 Hz AOM each
    C1 peaks

    C2 peaks

    C4 peaks



    ID11460 peak search with OD3 and 29ms probe
    49pts/peak. 150 Hz span, 3 Hz step
    Fourier limit is ca 25 Hz - so the 50/100 Hz sidebands are resolved.
    C1 peaks

    C2 peaks

    C4 peaks


    ID11458 peak search with OD1 and 2.9ms probe
    +C1 75 965 242, FWHM 174, h26
    -C1 75 959 314. FWH; 153, h34
    C1center 75 962 279, B1 1.0562 uT
    +C2 75 968 195, FWHM 159, h35
    -C2 75 956 362, FWHM 169, h31
    +C4 75 977 060, FWHM 152, h27
    -C4 75 947 479, FWHM 149, h18

    Estimated line-center from comb 75 962 255

    here get: A 6.7%, B 13.3%, C 8.4 %
    algorithm suggests [ 0.27459138 26.48031258 49.01349212]
    old voltages: 0.2693, 26.9714, 49.4475.
    ID11446 Try short uMotion-search 3x3x3 grid with 30kphotons
    uMotion 100kphot: A 15.3%, B 14.7%, C 9.6%

    15:56Z stop collition monitor
    cavity output-power now more than on Friday? ca 0.7 V
    comb f0 lock seems very unstable
    fiber was blocking ion-breadboard 674nm power-detector...
    turn on comb, comb-branch, clock laser
    15:50Z ion still stored

    2021-01-29

    16:27Z leave lab.
    16:20Z start collision-monitor with 10ms gate, 65/3 bright/dark. DB tag '2021-01-29-run1'. NO clearout.
    states
    dark_durations
    bright_durations

    Fluorescence levels 66 / 2.8 in 10ms
    Switch to B-beam.

    Turn off clock-laser and beat-note branch
    Comb f0 unlocked. (maybe lab-temperature changes when cavity-box was closed?)
    Cavity drifts too much to leave any useful clock-laser runs

    ID11439 re-start OD3/29ms search, with wider span, since we have so much uncompensated cavity-drift.
    71pts/peak
    C1 peaks

    C2 peaks

    C4 peaks


    ID11438 peak search with OD3 and 29ms probe (Fourier limit 13 Hz AOM)
    Note we have 50-60 mHz/s uncompensated cavity drift!
    C1 Gaussian fit fails...

    ID11437 repeat peak-search (since something jumped??)
    Cavity drift seems to be 70 mHz/s - since we touched the cavity when testing power-shift
    OD2, 9.9ms
    C1 75 964 506, FWHM 97 Hz, h24 and 75 958 555, FWHM 107Hz, h36
    C1center 75 958 555, B1 1.0604 uT
    C2 75 967 440, FWHM 107 Hz, h38 and 75 955 605, FWHM 123 Hz, h21
    C4 75 976 334, FWHM 94 Hz, h32 and 75 946 679, FWHM 89 Hz, h21
    duration 783s
    C1 peaks

    C2 peaks

    C4 peaks


    ID11436 peak search, OD2, 9.9ms probe (Fourier limit 40 Hz AOM). 8 Hz step, 170 Hz span
    C1 75 964 528, FWHM 93 Hz, h21 and 75 958 587, FWHM 103 Hz, h31
    C1center 75 961 557, B1 1.0587 uT
    C2 75 967 452 FWHM 112 Hz, h32, and jump??
    C4 only zeros?
    C1 peak

    C2 peaks

    C4 peaks


    ID11435 Repeat OD1 peak search for C1 pair. 3.3ms probe (Fourier limit 121 Hz AOM), 25 Hz step
    C1 FWHMs 134 Hz and 150 Hz
    C2 FWHMs 150 Hz and 164 Hz
    C4 FWHMs 133 Hz and 137 Hz
    (measurement time 688s)

    C1 peaks

    C2 peaks

    C4 peaks


    ID11422: Rabi flop with OD1, att. 18 dB. tau_pi 3.3 ms. -> use 10 ms for OD2


    ID11418 Line scan over C1+ with 18 dB att. and OD1.

    ID11415 Rabi flopping on C1+ with 18 db att. and OD1. Measurement was off-peak.

    ID11411 peak-search with OD1 and 3ms probe. 15 dB att. C4+ fit wrong. Corrected dataset manually.
    13:08 close cavity doors

    Try to measure cavity power-shift.
    ca12:20Z: 0.40 V power-servo set-point (servo measures cavity output-power with photodiode, and steers BOA 1348nm amplifier current)
    12:24 feedback off
    12:24:30 0.3V set-point, feedback ON
    1s comb-data noise is ca 25 Hz at IR frequency of 222.5 THz (1e-13, the maser noise level at 1s)
    5min = 300s average should give 1/sqrt(300) of this, or 1.4 Hz @ IR = 2.2e-14
    12:32 adjust power-servo, set-point 0.4V (servo on causes BOA current variations 450-490 mA)
    5min beat-note 23 161 465.41 (1s stdev 25.8)
    12:37 servo off
    12:39 servo set-point 0.3V (BOA current 350-380 mA) - servo-bumps on beat-note SA look much smaller
    5min beat-note 23 161 492.17 (1s stdev 27.2)
    12:44 servo off
    12:46 set-point 0.35V (BOA current 410-440 mA)
    5min beat 23 161 508.94 (1s noise 25 Hz)
    12:52 set-point 0.45V (490-530mA)
    in-loop photiode scope-signal at mu=451, 450.9, 451.0, 451.0, 450.9, 451.0, 451.2 451.1 451.0 451.1 mV
    5 min beat 23 161 522.1 (25.3 stdev)

    Power-shift slope is 190 Hz / V of frequency shift in IR, relative to cavity-output photodetector voltage
    This is 8.5e-13 fractional per volt.
    If we want a 1 Hz 674nm laser-stability, we need 0.5 Hz in IR, or 2e-15 fractionally
    This leads to a requirement of 2.6e-3 for the fractional voltage stability of the photodetector
    Closed-loop we see 0.3mV/450mV which is 7e-4 fractional


    uMotion contrasts with new voltages [ 0.26927269 26.97143157 49.44745276]: A: 2% in 11 s, B: 5% in 12 s, C: 2% in 17 s

    ID11403: NPL search for umotion minimum: 4x4x4. Result: [ 0.26927269 26.97143157 49.44745276] (optimum now within search-grid?)

    Cfit11403_inv =
    
       -0.0203    2.7627    2.2088
        0.3076   -1.4601    0.2457
        0.2695    1.1614   -3.8900
    

    Previous voltages: (0.3821,26.7839,51.0685)
    ID11400: NPL search for umotion minimum: 4x4x4. Result: [ 0.22398756 27.24888464 49.35180688]

    uMotion 100kphot: A 30% in 10s, B 46% in 11s, C 23% in 13s

    Peak search terminated (ID11392). Partial +C1 data in DB.

    2021-01-28

    (rabi-sinc with transform-limited FWHM shown for reference)


    ID11391, ID11392: Overnight: Peak search with OD2+OD3 filter, 225 ms probe
    DB-tag: 2021-01-28-search14-, -search15-.
    Fast drift comp. Estimated length of one search: 2.2 hrs per peak = ~13 Hr
    100 Hz span per peak. Step: 0.3 Hz, 333 pts/peak. Fourier limit 1.78 Hz AOM
    C1 peaks

    C2 peaks

    C4 peaks




    ID11387, ID11388: Rabi flop with OD1+OD3. tau_pi closer to 75 ms. -> use 225 ms probe for OD2+OD3.


    ID11386: Scan over +C1 with OD1+OD3 filters, 72 ms probe, DB-tag: 2021-01-28-run8. With fast drift-comp.


    ID11385: Peak search with OD1 filter, 3 ms probe, DB-tag: 2021-01-28-search12-. Fast drift comp.
    +C1: height 28, FWHM: 173 Hz
    -C1: height 42, FWHM: 178 Hz
    +C2: height 42, FWHM: 172 Hz
    -C2: height 27, FWHM: 147 Hz
    +C4: height 53, FWHM: 152 Hz
    -C4: height 30, FWHM: 163 Hz
    C1 peaks

    C2 peaks

    C4 peaks


    ID11382: Scan over +C1 with OD2+OD3 filters, 216 ms probe, DB-tag: 2021-01-28-run7. With fast drift-comp.


    ID11381: Scan over +C1 with OD1+OD3 filters, 72 ms probe, DB-tag: 2021-01-28-run6. With fast drift-comp.

    Disable drift-comp in crontab, and update drift-DDS every second: ID11378: Scan over +C1 with OD1+OD3 filters, 72 ms probe, DB-tag: 2021-01-28-run4. With fast drift-comp.


    ID11377: Scan over +C1 with OD1+OD3 filters, 72 ms probe, DB-tag: 2021-01-28-run3. With normal drift-comp.

    ID11376: Scan over +C1 with OD1+OD3 filters, 72 ms probe, DB-tag: 2021-01-28-run2. With normal drift-comp.


    ID11375: Scan over +C1 with OD2+OD3 filters, 216 ms probe, DB-tag: 2021-01-28-run1


    ID11373: Peak search with OD1+OD3 filters, 80 ms probe, DB-tag: 2021-01-28-search9-. Aborted after C1-

    ID11372: Peak search with OD1 filter, 3 ms probe, DB-tag: 2021-01-28-search8-. Data looks OK now.
    +C1: height 34, FWHM: 151 Hz
    -C1: height 47, FWHM: 162 Hz
    +C2: height 44, FWHM: 164 Hz
    -C2: height 27, FWHM: 151 Hz
    +C4: height 44, FWHM: 160 Hz
    -C4: height 26, FWHM: 170 Hz

    Switch back to A-beam
    It turns out, that C-beam was on?!?!

    ID11363: Rabi flopping without OD filters
    ID11351: Peak search with OD1, peaks seems really broad.

    re-lock comb f0
    uMotion 100kphot: A 18.2% in 10s, B 38% in 9s, C 9% in 18s

    08:40Z ion still stored. Clock-laser locked. Comb f0 unlocked sometime during the night
    60s stack of 1s exposures. old objective, ASI camera

    2021-01-27



    Scan results:
    ID11337 (run4), 11338 (run5), (18 local)
    ID11339, 11340 (19 local)
    ID11341, 11342 (20 local)
    ID11343 (21 local)
    ID11344, 11345 (22 local)
    ID11346 (run13), ca 23 local?

    overnight scans:
    OD3+OD1 filters, guess 72ms tau_pi (Fourier limit 5.6 Hz AOM). scan C1+ many times. 400 Hz span, 1Hz step
    each scan takes 4900 seconds = 1.4 hours
    10 scans in the queue, run3 thru run13


    ID11322 peak-search with OD2 and 7.2ms probe (Fourier limit 56 Hz AOM). 400Hz span, 10Hz step, 41 pts/peak
    C1+ h 24, FWHM 109 Hz
    C1- h 38, FWHM 99 Hz
    B1 1.0585 uT, linecenter 75 962 451
    C4+ h 32
    C4- h 18

    Try OD2 filter.
    ID11319 rabiflop on +C1, tau_pi 7.2 ms?


    ID11315 rabiflop on +C1.


    ID11314 peak search with OD3, 20ms probe (guess of tau_pi), Fourier limit 20 Hz AOM, 200 Hz span, 5 Hz step, 41 pts/peak
    +C1 75 965 411, h 13, FWHM 89 Hz but Gaussian fit does not fit central peak
    -C1 75 959 471, h 21, FWHM 76 Hz. Central peak ca 25 Hz AOM wide
    B1 1.058 uT
    linecenter 75 962 441 Hz
    C2+ 75 968 369, h 22, FWHM 96 Hz
    C2- 75 956 527, h 18, FWHM 79 Hz
    C4+ peak is far left in the span, B-field estimate too high?
    C4- peak is far right in span

    Repeat OD1 search (ID11313), Influx tag: 2021-01-27-search6-
    all peaks roughly transform-limited with 3 ms probe


    Rerunning peak search with OD1 (ID11312), probe 3 ms, Fourier limit 133 Hz AOM, Influx tag: 2021-01-27-search6-: C1+ heigth 33 FWHM 154 Hz
    C1- height 40 FWHM 166 Hz
    C2+ height 34 FWHM 191 Hz
    C2- height 24 FWHM 166 Hz
    C4+ height 49 FWHM 155 Hz
    C4- height 29 FWHM 152 Hz

    Problem with saving Zeeman-peak data in peak-search program. Problem corrected.

    Try OD3 filter, should work with 27ms?
    ID11286 peak search, A-beam, OD3 filter, 27ms probe. Fourier limit 15 Hz AOM. 150 Hz span, 3 Hz step, 49 pts/peak
    (db tag 2021-01-27-search5-)
    +C1 central peak looks OKish, but Gaussian fit is to all of data. no clear 100Hz sidebands?
    Other peaks looks similar
    Clock probe maybe too long?

    Try OD2 filter, should work with 9 ms probe
    ID11285 peak search, A-beam, OD2 filter, 9ms probe (Fourier limit 44 Hz AOM). 250 Hz span, 10 Hz step, 25 pts/peak
    (Note Fourier limit close to AC frequency of 50Hz - could interfere?)
    (db tag 2021-01-27-search4-) +C1 h 22, FWHM 136 Hz AOM - looks noisy, Gaussian not a great fit (span is a bit narrow)
    -C1 h 34, FWHM 115 Hz - B1 1.065 uT - linecenter 75 962 441 Hz
    +C2 h 33, FWHM 140 Hz - strong sidebands at 5 steps = 50 Hz AOM = 100 Hz optical from center! - Gaussian is fit to everything, not only central peak
    -C2 h 21, FWHM 153 Hz - also sidebands
    +C4 h 37, FWHM 104 Hz - highest peak in this run! - not any wider than C1/2, so isn't limited by B-field?
    -C4 h 79, FWHM 110 Hz - Gaussian covers central peak and side-bands..
    (duration 903s)

    Try OD1 filter, this should work with 3ms probe
    ID11283 (data file stored as ID11284?) peak search, A-beam, OD1 filter, 3ms probe (Fourier limit 133 Hz AOM). 500 Hz span, 25 Hz step, 21 pts/peak
    (db tag 2021-01-27-search3-)
    +C1 h 33, FWHM 157 Hz AOM
    -C1 h 46, FWHM 157 Hz - B1 1.066 uT - linecenter 75 962 434 Hz
    +C2 h 45, FWHM 142 Hz
    -C2 h 25, FWHM 175 Hz
    +C4 h 45, FWHM 164 Hz
    -C4 h 26, FWHM 183 Hz
    (duration 683s)


    ID11275 NPL uMotion search over lunch. 4x4x4 grid
    algorithm suggests Optimal voltages (Common mode, diff): [ 0.38212489 26.78389499 51.06850713] - both first and last value are outside scan range.
    Try suggested voltages.
    uMotion 100kphot: A 4.7% in 14s, B 27.3% in 12s, C 12.9% in 14s. (overall better than before, but not optimal)

    ID11274 peak-search with 1000us probe (Fourier limit 400 Hz AOM). 1.5kHz span, 80 Hz step, 19 pts/peak
    +C1 height 43, FWHM 384 Hz
    -C1 h 52, FWHM 423 Hz - B1 1.064 uT - liencenter 75 962 455 Hz
    +C2 h 50, FWHM 405 Hz
    -C2 h 41, FWHM 356 Hz
    +C4 h 51, FWHM 432 Hz
    -C4 h 30. FWHM 415 Hz
    (duration 607 seconds)

    ID11272 rabilfop on +C1, to 5 ms probe. tau_pi ~ 1ms (this might give better peak-heights than previous peak-search)

    ID11271 rabiflop on +C1

    Peak search (ID11270, 18 dB att., 500 us probe):
    +C1: Center = 75.9654 MHz, height 20, FWHM: 890 Hz.
    -C1: Center = 75.9594 MHz, height 27, FWHM: 942 Hz.
    +C2: Center = 75.9684 MHz, height 35, FWHM: 873 Hz.
    -C2: Center = 75.9565 MHz, height 22, FWHM: 861 Hz.
    +C4: Center = 75.9773 MHz, height 22, FWHM: 862 Hz.
    -C4: Center = 75.9476 MHz, height 12, FWHM: 840 Hz.
    B estimate = 1.07137 uT.

    Coil currents set to (-8,0,0).
    -3 floor A/C setting is 20 Hz

    NOTE uMotion contrasts not optimal, all beams ca 30% now!

    ID11260 compare to non-zero B-filed by flipping I1 sign, currents: (+4.4719, -1.0115, 1.0309) mA. lower 18 kcounts/s at max
    ID11258 repeat A-beam detuning scan, 1000ms gate. Far detuned 2 kcounts/s, near resonance max 21.3 kcounts/s
    ID11257 A-beam detuning scan - no Dork-resonances apparent..

    ID11254 clock-spectrum with zero B-field, 500us probe (Fourier limit 800 Hz AOM), 3 kHz span, 50Hz step
    Gaussian fit FWHM 908 Hz
    center 75 962 438
    height 88

    ID11252 rabiflop on center, in zero B-field, visually tau_pi ca 500us

    ID11248 clock-spectrum with zero B-field?
    A-beam, 600us probe (Fourier limit 667 Hz AOM), 100Hz step, 5 kHz range. zeeman AOM attenuation 18dB from y-day
    Gaussian fit FWHM 760 Hz AOM
    center 75 962 481 Hz


    Try the zero B-field currents I1 = -4.4719, I2 = -1.0115, I3 = 1.0309

    uMotion 100kphot: A 36.3% in 9s, 30.4% in 10s, 28.7% in 13s

    08:19Z ion still stored, clock-laser locked to cavity, but comb f0 was unlocked.

    Initial coil results:
    B = B0 + A I
    
    Ab =
        0.2649(2)         0         0
             0    0.41962(14)         0
             0         0    0.2713(2)
     
    B0b =
        1.1847(14)
        0.4244(9)
       -0.2796(14)
    % current giving zero field
    IO = [-4.4719   -1.0115    1.0309] mA
    

    2021-01-26

    Overnight coil-characterization
    A-beam, 600us probe, 500Hz step (note this is quite sparse sampling... Fourier limit 667 Hz AOM), 50 kHz span, 101 points per spectrum
    ca 5s per point, so only 500s per spectrum. We have 3*6 spectra, so 3*6*500 = done in 3 hours.
    step each current: +2, -2, +4, -4, +8, -8 mA.
    2021-01-26-coil-0-0-0 is the first run, with zero currents
    0, 0, 0
    +2, 0, 0
    -2, 0, 0
    +4, 0, 0 leave lab around 17:10 local, artiq got to this point.
    -4, 0, 0
    +8, 0, 0
    -8, 0, 0

    0, +2, 0
    0, -2, 0
    0, +4, 0
    0, -4, 0
    0, +8, 0
    0, -8, 0

    0, 0, +2
    0, 0, -2
    0, 0, +4
    0, 0, -4
    0, 0, +8
    0, 0, -8

    Note the last coil characterization results from 2020-08-19: (coil sensitivities very similar on 2020-06-12)
    % 2020-08-19
    A2 = [0.216895938211581                   0                   0;
                       0   0.325633046718816                   0;
                       0                   0   0.208504965524669];
    B02 = [1.085916150083249, 1.291554002640864, -0.100996485194311]';
    

    orange: zero coils
    blue: I1 varied
    red: I2 varied
    green: I3 varied


    Remove OD-filters, for overninght coil characterization
    A-beam, no OD filters, rabiflop shows 600us tau_pi


    Spectrum with 4.5ms and 18dB attenuation. Fourier limit 89 Hz AOM
    Gaussian fit FWHM 124 Hz AOM
    center 75 963 825 Hz

    Try increasing Zeeman AOM attenuation, to only slightly lower probe intensity. Attenuation 18dB.
    Rabi-flop on center gives 4.5ms tau_pi


    ID11197 scan with OD2 and 3.2ms. Fourier limit 125 Hz AOM
    Gaussian fit FWHM 144 Hz AOM
    Center 75 963 826
    ID11195 rabiflop on center, OD2 filter. tau_pi 3.2ms

    OD3 filter is too hard? go back and try OD2 filter?

    re-run rabiflop on center

    re-run +C1 scan with 9400us probe (Fourier limit 42 Hz AOM). A-beam, OD3 filter
    Gaussian fit FWHM 102 Hz AOM

    Rabiflop, to find tau_pi?

    OD3 filter, 15ms probe, scan around +C1


    OD2 filter, 4500us probe, scan around +C1


    Try coil currents (-4, 0, 0) mA, scan with 450us


    Peak-search with A-beam and 450us probe:
    but doesn't find C1.

    rabiflop on -C1, visually tau_pi ca 450 us
    P674_A = 0.8 uW, no OD-filters (note B-beam had about 5uW!)
    Try A-beam, it may have better 422nm polarization?
    ID11174 scan with 600us probe pulse


    C1 peak search OD3+OD1 filter, 15 ms probe, 5 Hz step, 100 Hz span, Fourier limit 27 Hz AOM
    +C1 not a great peak..
    -C1 75.960007, FWHM 70 Hz AOM, h 22 (Gaussian fit not so great)

    C1 peak search, OD3 filter, 5400us probe pulse, 300 Hz span, 20 Hz step, Fourier limit 74 Hz AOM
    +C1 75.964675, 129 Hz, h 19
    -C1 75.960007, 124 Hz, h 35
    B1 = 0.832 uT

    C1 peak search, OD2 filter, 1800us probe pulse, 500 Hz span, 50 Hz step. Fourier limit 222 Hz AOM
    +C1 75.964672, 272 Hz, h 19
    -C1 75,962336, 246 Hz, h 48
    B1 = 0.832 uT

    C1 peak search +C1 75.964650, 703 Hz AOM FWHM, height 30
    -C1 75.962345, 753 Hz FWHM, height 32
    B1 0.821 uT

    Coil sensitivity roughly dB/dI1 = 0.24 uT/mA ?? (compare to previous results)
    abs(B0) field at zero coil currents 1.3 uT?
    Lineshape scan, OD1: ID11132
    Change coil currents to (-2,0,0)


    Lineshape scan, OD1: ID11130. C2 peaks (?): 75.97232, 75.95232
    Change coil currents to (2,0,0).

    Install the OD1 filter, this should increase tau_pi 3-fold, to 600us.
    ID11121 scan spectrum with OD1 filter, 600us probe, 40 kHz range, 200 Hz step. Duration 800s


    Comb f0 had unlocked over lunch, fixed by adjusting waveplates in GUI

    All peaks seem to fit within 40 kHz AOM
    ID11120 wider spectrum scan with 200us probe
    C1 separation looks like roughly 15 kHz AOM ?


    ID11119 rabiflop on +C1(?), tau_pi ca 200us
    ID11118 initial clock transition scan with 15dB AOM attenuation, 600us probe pulse.

    674 power in B-beam: 5 uW

    Detuning scans: A: ID11108, B: ID11110 C: ID11112. Gate 20 ms. Counts: ID11115 (dark approx. 10)

    Manually trying: (0.28, 27.3333, 53.3333): 8%, 6%, 3%

    Automatic NPL search (ID11098) gave weird result [ 0.08607781 32.11213181 52.31081692]

    Try (0.3571, 26.8571, 53.5714), got 3%, 31%, 3%

    -0.05B step to (0.1334, 26.77556, 54.59336) gives (3%, 2%, 39%)
    Repeating search with (0.2489, 26.6999, 54.6666) gives now 7%, 25%, 29% (repeat 10%, 29%, 36%)

    NPL search finished with [ 0.33088186 27.33750859 54.74849937], but ion seems to have disappeared during the night.
    Last measured contrast seems ok.
    Ion reappeared after switching on all beams.
    Contrast with new voltages: 33%, 32%, 16%

    2021-01-25

    Leave NPL-algorithm running, 8x8x8 grid
    ID 110100

    By scanning V34diff, found B and C optimal point:
    Contrasts: 30 1.4 3.9 %
    Voltages: 0.2489, 26.6999, 54.6666

    Found rough optimum for B beam, by stepping along the B-vector
    Contrasts: 71, 13, 69 % at voltage of (0.841, 26.692, 47.1523)

    From 2020-10-05 we have some notes on the cardinal directions:
    Cfit9327_inv =
        0.2372    2.3100    1.9737
        0.4131   -1.5132   -0.0264
        0.6644    1.4648   -3.8875
    
    uMotion 100kphot: A 70% in 10s (B and C similar ca 70%)

    Coil currents from December were (127.7, 6.1, 18.4) mA. Now set them all to zero

    Adjust alignment into 844nm amplifier - seems to improve 422nm power
    422nm power seems halved compared to start of day - fiber coupler between 422-laser and ion-breadboard?
    uMotion 100kphot: A 73% in 58s

    Install mid shield
    Install inner shield
    Install pt100 sensor on IR side window flange

    uMotion 100kphot: A 82 % (anharmonic) in 38s, B 73% in 56s, C 76% in 75s (Note no shields!)

    C ca 70 counts/50ms
    B ca 100 counts/50ms
    A-beam and ASE alignment optimization gives ca 140 counts in 50 ms
    Found PMT signal, in upper left corner of camera image, as before.
    ion is still stored, Z-focus of camera was wrong..

    Align PMT-splitter with help of 422-fiber-end placed at PMT position. This allows aligning the PMT-splitter so 422nm-light is bright at analog camera on IR side
    Lost ion (?) while swapping cameras etc.
    Install the PMT-splitter next

    Switch to the ASI camera. Remove intensifier/PCO.

    Align B-beam (thru-transmitted diffraction pattern) -> get fluorescence (slightly weaker) with only B also.
    Align C-beam -> we get fluorescence with only C now also.
    Try camera (z) focusing - but it was already quite optimal. (thus aligning the camera to the electrodes, and defocusing by 750um/2 is a good way to z-focus the camera)
    set detuning to 208 MHz - more fluorescence
    optimize A and ASE directions, but no major fluorescence changes. PCO camera to 200ms exposure
    11:53:15 see single(?) ion? load time 5min 15s
    11:48 try to load single ion. 1.7A dispenser


    throw away cloud by shutting off RF
    A cooling shows cloud, B and C quite dark (do not cool much?)
    11:43:37 dispenser off. cloud remains
    11:42 see a cloud! (intensifier gain higher, PCO exposure up to 1.2 seconds). A-beam adjustments while dispenser on finally revealed the cloud.
    stopping down the imaging objective with the iris reduces a vertical strange 422-reflection in the camera image
    11:27:15Z try load, dispenser 1.7A. Intensifier gain 2.7V, PCO camera 200ms exposure

    Align ASE beam with the analog camera, A-beam with the thru-transmitted diffraction pattern from the electrodes.
    Optical power in the beams:
    422nm A 0.9 uW, B 1.1uW, C 1 uW. 185 MHz detuning
    461nm A 180 uW
    clearout 65 uW
    repumper 10.8 mW
    405nm A 3.6 mW

    RF 14.424 MHz, 10Vpp generator, 2.759 Vpp monitor (scope in averaging mode)

    Latest bias-voltages are from 2020-10-29 (?). They are -6.2, 0.2635, 27.0703, 46.7861. Apply these.
    This gives individual voltages: V1 = -6.4635, V2 = -5.9365000000000006, V3 = -19.715799999999998, V4 = 73.8564

    Analog camera images IR/ASE side. Intensifier + PCO images ion.
    Remove mid top-shield, and inner shields - to allow aligning beams. Mid bottom-shield(s) need not be removed?
    Try to trap again?

    2021-01-24

    Temperature rise with thermal compound between helical-resonator and alu-mount.
    Adding thermal compound changed helres heating from 120 mK to 90 mK.

    2021-01-22

    Installed longer antenna: monitor voltage now 2.78Vpp! f = 14.422 MHz

    re-tune SWR close to 1.0. Monitor 60 mVpp with 10Vpp drive from siggen
    Install heatsink-compound between helical-resonator and alu-mount

    Install inner and mid-shields. Seems to fit OK. Beams go through shield-holes OK.
    Note back feet of mid-shield bottom left/right half did not fit the designed M3 holes - they loose between bottom shields and pump-chamber

    Tickler voltages measured after coupling capacitor (at eletrode) with 10x probe. Voltages are peak-to-peak

    Measured voltages seem low. Capacitors of the DC filter (fig. from OC18 guidelines) equal to approx. 0.5 or 0.25 ohms at 1 MHz and 2 MHz respectively.
    This seems to load the DDS output significantly

    Without attenuator

    f (MHz), DDS att., V

    1.0 5 distorted..
    1.0 10 7.4 mV
    1.0 15 4.4 mV
    1.0 20 2.9 mV
    2.0 5 distorted..
    2.0 10 3.0
    2.0 15 2.0 mV
    2.0 20 1.5 mV

    20 dB attenuator

    f (MHz), DDS att., V

    1.0 3 1.97 mV
    2.0 3 ~500 uV (noisy)

    2021-01-21

    Mu-Shields delivered. Inner and mid-shield unpacked - looks ok. Quite heavy duty material thickness!

    Tomorrow: Check ASE-beam alignment (analog camera), RF trap drive OK?
    15:03 no signal, dispenser off
    Intensifier gain turned up, PCO exposure ca 100-200ms
    14:51Z try to load with 2.0A. Now using image-intensifier and PCO camera.

    Field of view now larger, both top and bottom almost in same view
    NOTE: Image-intensifier input side is closer to the electrical wires. 3 wires: Orange (+5 to +15 V supply, 30 mA), White (Control voltage 0-5 V), Grey (Ground).
    Use the old PCO-camera, to image the intensifier output plane
    Install the (new) image-intensifier, this may help detect weak signals?

    12:56 no signal, dispenser off
    12:46 try load with 2.0A dispenser (2.3W initially)

    check Rb-lines, they look ok
    check 461nm with dispenser: 198.2 mA, 4.645 kOhm
    12:32Z no signal, dispenser off
    12:17Z try to load, 1.7A dispenser
    remove PMT-splitter, to get more signal to camera. Use old imaging-objective. re-align objective to electrodes
    (O-ring inside PMT-splitter-cube was loose?) blocking light thru lid is important

    11:07Z no signal, dispenser off
    10:53Z try to load, 1.7A dispenser

    09:46 no signal, dispenser off.
    09:32Z try to load, 1.7A dispenser
    with old objective A-beam window reflection is quite strong, and exits the objective-tube towards B-beam post
    new objective may have issues with beam window-reflections? Brass tubes used as spacers may need to be painted black?
    test new imaging objective - but no huge improvement in image-quality of electrode image with flashlight illumination
    realign camera to electrode image

    08:52 no signal, dispenser OFF
    08:37Z load with 1.7A dispenser
    trap RF as before. monitor waveform seems jumpy!?

    Optical power (slim sensor)
    422nm A-beam 0.8uW, B 1.1 uW, C ~1uW
    461nm A 200uW
    405nm adjust to 3.7 mW
    clearout 64 uW
    repumper 10.6 mW

    align A-beam with help of electrode reflections
    remove inner shield (outer shield was removed yesterday).

    win machine tilted? reboot.

    2021-01-20


    13:28 no signal, dispenser OFF
    adjust A-beam, based on output beam shape...(??)
    13:15 try 1.7A dispenser, 1.7W
    13:13 no signal, dispenser OFF
    13:01Z try again with 1.6A Dispenser. 422nm Rb-lines checked, 405nm power checked. A-beam output doesn't look very clean (hits some aperture somewhere?)

    12:50 no signal, dispenser OFF
    12:47 no signal in 10min..
    12:37:30Z dispenser on, 1.6A. 1.45W initially, rising slowly to 1.5 W
    Try to load with 1.6A dispenser current

    Trap RF from signal generator 10 Vpp, 14.424 MHz
    monitor 64 mVpp

    Preparing for trapping.
    461 dispenser check, OK at I = 198.3 mA, T = 4.645 kOhm.
    (New Thorlabs slim power sensor, model S130C)
    422 power: A exit: 0.85 uW, B: 1.1 uW, C: 1.0 uW. Detuning at 185 MHz
    461nm A-beam 195 uW
    405nm A-beam adjust to 3 mW
    repumper 13.5 mW (repeat 12mW) at ASE beam output
    clearout 65 uW (seems low?)
    (AW paper notebook clearout values: 75 uW on 2019-12-04, 77uW on 2019-11-25, 78uW on 2019-11-04)

    New Kubuntu linux install

    2021-01-19

    Installed new OS on new disk. Migrated all software.
    Artiq version now: v5.7144.9ed8f663. Conda enviroment: artiq-ion-v5-6.
    Ionclock ARTIQ now ident 5.7144.9ed8f663;mikes (autogenerated by M-Labs).
    New drift-DDS ARTIQ now ident 5.7144.9ed8f663;mikes-drift (built manually).
    Drift-DDS ARTIQ seems to work. Configured for external ref for both Kasli and Urukul.

    2021-01-16

    (slight improvement with lid-adjustment and heatsink)

    2021-01-14 Thursday


    13:50Z add a heatsink to the alu end-plate
    13:34Z improve slightly contact between antenna-lid and end-alumount. No visible reduction in monitor-voltage when loosening lid M3 screws (impedance matching still ok?)

    With added clamping, resonator heating is reduced 140 -> 120 mK. RF-feedthru heating ~few mK lower.
    (note the new year 2021 AD...)

    2021-01-13 Wednesday

    Install 2pcs hose-clamps that push the Cu-resonator into contact with the Alu-mounts
    12:13Z install clamps to breadboard on both sides of Alu helres-mounts

    Heating of the RF-feedthrough flange (10Vpp from signal-gen, 6+ hours settling time):
    Old mount, no sink 120 mK
    Old mount, Cu bars as heat-sink 85 mK
    New mount, 2 mounts clamped 66 mK
    New mount, mounts clamped, 2 hose-clamps around cu-helres 66 - ~few mK (results 2021-01-14!)

    Results with two alu-mounts clamped to breadboard, and no clamping-force between cu-resonator and alu-mounts.

    2021-01-11 Monday

    Install new alu-mounts that imrpvoe heat-ocnduction from Cu-resonator to breadboard

    11:26Z remove temporary copper heatsinks from helres

    2021-01-08

    (TLi) New vented rf shield (0.1 mm copper) with 6-mm slots top and bottom:
    ca 12Z move Pt100 ch. 6 (106) from rf shield to helical resonator cylinder, above adapter piece.

    With old shield:
    Signal generator (Vpp) Monitor (mVpp, average 32)
    10 66.4 (without averaging ca 70)
    9 59.5
    8 53.5
    7 47.0
    6 40.2

    rf off.
    New shield installed. RigExpert: min SWR 1.02 @ 14.4245 MHz, R = 52.1 Ohm, Q ~ 870 (SWR 2) ... 810 (SWR 3).

    rf on.
    Signal generator (Vpp) Monitor (mVpp, average 32)
    8 53.1
    10 66.8

    Both Q and monitor ratio as before within uncertianty.

    ca 13Z ready

    2021-01-07

    14:15Z RF OFF. will cool down by midnight, when computer turns ON.

    RF siggen to 10Vpp, turn ON for testing
    Menlo comb
    lab lights (partially)
    ca 09Z Turn on stuff after x-mas break

    For 2020 results see index_2020

    For 2019 results see index_2019